### v1

# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

Abstract
Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices. We introduce MetaLint, a new instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications. Unlike conventional approaches that train models on static, rule-based data, MetaLint employs instruction tuning on synthetic linter-generated data to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining. To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize. Our results show that MetaLint improves generalization to unseen PEP idioms, achieving a 70.37% F-score on idiom detection with the highest recall (70.43%) among all evaluated models. It also achieves 26.73% on localization, competitive for its 4B parameter size and comparable to larger state-of-the-art models like o3-mini, highlighting its potential for future-proof code quality analysis.
1 Introduction
With the rise of Large Language Model (LLM) agents, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent. Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023). However, many of these efforts are hindered by reliance on task-specific training data, often based on narrow or outdated coding best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024). As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024). They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). Ideally, we would develop LLM systems capable of identifying code quality issues without explicit supervision for the target idioms, especially for hard or rare patterns, and that can adapt to evolving best practices over time.
One approach to this problem is to train the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms. For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function. However, as noted in PEP 506 (D’Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against the usage of this module for security reasons, which is often missed by developers as highlighted by accepted answers on forums like StackOverflow. PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom. As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024). Additionally, the issues can be fixed by replacing instances of “bad” idioms with corresponding “good” idioms that align with best practices.
In this work we extended this observation by training the LLM on a higher-level instruction-following task: given a specification of a best-practice code idiom $I$, the model learns to identify and localize code fragments that violate the idiom and could be rewritten to adhere to or incorporate idiom $I$. We refer to this instruction-following task as meta-linting. The core idea is to teach the LLM the underlying structure of the task, namely identifying and locating idioms based on high-level specifications, rather than training it to memorize or learn a static set of best practices. Since the focus is on improving the model’s instruction-following ability rather than tailoring it to a particular data distribution of best practices, this approach can keep up with evolving best practices.
Our approach, MetaLint, is motivated by the observation that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Since meta linting frames each idiom as a separate task, instruction style training is naturally suited to generalize to novel and unseen idioms. Existing linters for programming languages (e.g., Ruff for Python) implement numerous simple rules, or “easy” idioms, that reflect aspects of code quality such as modern syntax, readability, and security. These linters can be used to generate large-scale synthetic data for supervised instruction fine-tuning and serve as verifiers during preference optimization or reinforcement learning. This setup also supports easy-to-hard generalization: training on automatically generated linter data for easy idioms helps build a strong model that can handle harder, more abstract idioms at test time, which go beyond the reach of static analyzers. Similar strategies have shown success in math reasoning tasks (Sun et al., 2024b). To systematically study this generalization, we construct a challenging benchmark of hard idioms based on popular PEPs that introduce high-level constructs. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them to models trained with MetaLint, focusing on whether LLMs can move beyond memorization, generalize from easy to hard idioms, and adapt to novel patterns across code quality domains.
Our key contributions are as follows:
  1. 1.
    We introduce MetaLint, an instruction-following framework that leverages synthetic data for easy-to-hard generalization and addresses the limitations of static training used in prior LLM-based code quality approaches.
  2. 2.
    We construct a benchmark of challenging, widely relevant idioms inspired by various PEPs, and evaluate state-of-the-art code and reasoning models on it.
  3. 3.
    We analyze whether MetaLint promotes adaptive reasoning over memorization and enables generalization from easy to hard idioms.
2 Related Work
2.1 Code Quality Analysis with Large Language Models
A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis. Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as linters guided by rule descriptions, relying on prompting or fine-tuning to improve rule enforcement. LiteLLM (Fang et al., 2025) leverages LLMs to reduce both the cost and false positive rate of linting for Verilog. Khare et al. (2023) show that LLMs can outperform traditional static analysis tools for detecting certain security-related CWEs, especially when step-by-step reasoning is used. Vijayvergiya et al. (2024) train LLMs to detect and localize violations of best practices, while Rasheed et al. Rasheed et al. (2024) design a multi-agent review pipeline targeting issues like maintainability, efficiency, and bugs. Other works (Jiang et al., 2025b; Yao et al., 2025) use comparative prefix-tuning and reinforcement learning with static analysis–based rewards to promote the generation of higher-quality functionally correct code. Naik et al. (2024) and Jaoua et al. (2025) integrate LLMs with linters to produce more informative code reviews. Finally, CoUpJava (Jiang et al., 2025a) presents a benchmark of Java version–related code upgrade histories, conceptually similar to the hard PEP idiom benchmark we developed for Python. While such efforts highlight the potential of LLMs in code quality tasks, they often target specific rule sets or best practices, necessitating retraining as those evolve. In contrast, our instruction-following meta-linting framework trains models to interpret high-level code quality specifications and perform static analysis to enforce them, enabling broader generalization beyond fixed rules.
2.2 Instruction Following for Generalization
Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning a single fixed task. Prior work shows that exposing models to diverse task instructions allows them to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, acting as a substitute for supervision (Puri et al., 2022) and enabling models to generalize even with minimal labeled data or via pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity is a key driver of generalization, with varied instructions improving performance more than repeated exposure to the same task format (Charton et al., 2024). This phenomenon holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) also performs task-specific instruction following using synthetic data, demonstrating its effectiveness; however, it relies entirely on LLM-generated data and does not incorporate verifiers. Collectively, these results suggest that instruction tuning functions as task-level meta-learning, equipping models with the ability to rapidly adapt to new tasks via natural language. Inspired by this work we treat modeling of specific code quality idioms as individual tasks and propose the MetaLint training framework that generates synthetic data at scale for each meta-task to facilitate cross-idiom generalization. This way the trained model can keep up with new idioms and evolving best practices.
2.3 Easy-to-Hard Generalization
Recent research demonstrates that easy-to-hard generalization improves model performance across mathematical, algorithmic, and coding tasks, motivating its application to code quality analysis. In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025). Several works implement curriculum strategies to structure this progression (Parashar et al., 2025; Zhao et al., 2025), while others emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024). These principles extend beyond math: Sun et al. (2024a) shows that reward models trained on simple code and math problems improve performance on complex ones, and Zhao et al. (2025) further explores generalization in code tasks through curriculum fine-tuning. Multi-task and length generalization studies (Hu et al., 2025) and differentiable programming setups (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code. Zhang et al. (2024) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats. While our approach does not explicitly follow curriculum learning, we draw inspiration from this line of work by training MetaLint on large-scale synthetic data that captures easily detectable code idioms, many of which are directly covered by rule-based linters. We hypothesize that these simple patterns act as a stepping stone, enabling generalization to more complex and abstract unseen PEP idioms.
3 Method
Figure 1: MetaLint training framework (1/2): Core components of our framework include using linter tools (like Ruff) for 1) Synthetic Data Generation, 2) Supervised Instruction Fine-Tuning (SFT) with synthetic data and a 3) Verifiable Reward Model (also derived from the Ruff linter).
Figure 2: MetaLint training framework (2/2): We use the linter-based verifiable reward model for the final two preference optimization components: 4) Rejection Sampling Direct Preference Optimization (RS-DPO), and, 5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
We propose the MetaLint framework to support the learning of skills essential for code idiom analysis, such as interpreting idiom descriptions in terms of underlying code constructs and performing line-by-line static analysis to detect idiom violations. We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}={D_{I},E_{I}}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
3.1 Synthetic Data Generation
A key advantage of our meta-task formulation is its support for easy-to-hard generalization. We train LLMs on a set of idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$ to improve their ability to flag idiom violations, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by $\mathcal{L}$). Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand, and detect code idioms from high level descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to an untrained checkpoint.
Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For this purpose, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, semantics, and more. We run Ruff on Python source code files $f\in\mathcal{F}$ from the STACK-V2 (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, the Ruff website hosts rule-specific documentation, including descriptions and examples, which we scrape to construct meta-task prompts $M_{I_{\mathcal{L}}}$ for each idiom. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix A.1.1.
3.2 Instruction Supervised Fine-Tuning
As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of 50 linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model’s output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix A.1.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some idioms, the final distribution is approximately 70:30 in favor of files with no violations. This results in a total of 53k training instances spanning 50 idioms.
3.3 Verifiable Reward Model and Preference Optimization
For preference optimization, we adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model. It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2). We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter’s line numbers (yellow circle in “Verifiable Reward Model”, Figure 1) as ground truth and the model’s predicted lines (blue circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$.
Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
$\Phi^{RL}=\arg\max\sum_{(x,y_{win},y_{lose})\in\mathcal{D}_{p}}\log\sigma\left(\beta\log\frac{\Phi^{RL}(y_{win}|x)}{\Phi^{SFT}(y_{win}|x)}-\beta\log\frac{\Phi^{RL}(y_{lose}|x)}{\Phi^{SFT}(y_{lose}|x)}\right)$
Here, $\sigma$ denotes the sigmoid function, and $\beta=0.1$ is the KL penalty coefficient, corresponding to low-to-moderate regularization.
3.4 Training with Reasoning Traces
Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning. To obtain CoT traces that help the LLM arrive at the correct answer, we adopt a rejection sampling approach (“Rejection Sampling SFT” in Figure 2) for collecting SFT data. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained LLM capable of CoT reasoning (e.g., Qwen3-4B), and compute the reward $r_{y_{i}}$ for each response, following a similar procedure as in RS-DPO. However, instead of forming contrastive pairs, we discard any $y_{i}$ with reward $r_{y_{i}}<\gamma$, where $\gamma=1$. In other words, we discard CoT–response pairs that are incorrect or improperly formatted and retain only those that are fully correct. The reward is applied only to the final response, obtained after extracting or parsing the CoT trace. We also remove any cases where the CoT fails to terminate or produce a final answer.
For some inputs $x$, no valid $y_{i}$ may be found. In such cases, we skip the input entirely. To promote meta-task diversity, we retain at most two valid responses $y_{i}$ per input. We sample multiple $y_{i}$ only for inputs with at least one violation, and only a single $y_{i}$ for inputs with no violations, this helps maintain a similar no-violation-to-violation ratio of 71:29 as in the regular SFT data. The latter case is also more likely to yield a valid CoT–response pair within the token limit. When more valid responses are available than needed, we keep the shortest $y_{i}$ completions, as they typically reflect more concise reasoning (the final answer usually consumes a similar number of tokens across samples). Following this policy, we collect 52.7k instances, which we use to train the base model using SFT.
After training the LLM on this dataset, we obtain a CoT-capable SFT model $\Phi_{CoT}^{SFT}$. We then apply the same RS-DPO procedure described in Section 3.3, with the only modification being that each $y_{i}$ now contains both the CoT trace and the final response.
4 Experiments
4.1 PEP Hard Idiom Benchmark
To evaluate whether MetaLint training helps the base LLM learn to interpret high-level idiom specifications, and apply this skill to harder, more abstract idioms that linters struggle to detect, we construct a benchmark of “hard idioms” sourced from PEPs. We begin by identifying fifteen PEPs that define semantic or abstract behaviors beyond what syntax or token patterns capture, then design specific heuristics for each to detect code that violates their guidelines. Using these heuristics, we search the STACK‑V2 corpus for violations, prioritizing high recall over strict precision to retrieve a broad set of candidates for manual selection. These “hard idioms” cannot be reliably detected by simple pattern matching, making them ideal for testing whether models truly understand idiom specifications or memorize frequent patterns seen in the Ruff idioms. From the retrieved examples, we handpick 15–20 representative files per PEP that clearly illustrate each violation. For each file, we identify the exact lines showing the violation (“before” snippet) and author a corrected version (“after” snippet) that complies with the PEP, forming the ground truth. To balance the benchmark, we include negative examples (code without violations) by selecting files retrieved for one PEP and confirming they don’t violate others. This results in 536 challenging examples, with 52% containing at least one violation and 48% being violation-free, allowing us to evaluate MetaLint’s ability to generalize from easy to hard idioms.
4.2 Evaluation Metrics
We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs. For both tasks, we compute precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each idiom, treating each idiom as a distinct class. Since the distribution of idiom occurrences may be imbalanced, we use macro averaging across idioms to weigh each class equally.
Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each instance $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$, we define the indicator variable:
$\mathds{1}[x]=\begin{cases}1&\text{if }x\text{ is true}\\ 0&\text{otherwise}\end{cases}$
The idiom detection metrics per idiom $I$ are defined as:
$P_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|=0]\cdot\mathds{1}[|\hat{y}|>0]\right)}$
$R_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|=0]\right)}$
We compute the final detection metrics by macro averaging across all idioms:
$P_{\text{Det}}=\frac{1}{|I|}\sum_{I}P_{I},\quad R_{\text{Det}}=\frac{1}{|I|}\sum_{I}R_{I},\quad F_{\text{Det}}=\frac{2P_{\text{Det}}R_{\text{Det}}}{P_{\text{Det}}+R_{\text{Det}}}$
For idiom localization, we compute set level precision, recall, and F1-score for each instance $(x,y)\in\mathcal{D}$, comparing the predicted and ground truth violation lines:
$P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
4.3 Baselines
We evaluate a range of state-of-the-art open and closed-source LLMs, trained for coding and reasoning tasks, on the PEP hard idiom benchmark to contextualize the improvements achieved by the MetaLint framework in flagging idiom violations from high-level specifications. Among open source models, we include instruction-tuned variants from the Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), and Qwen3 (Team, 2025) families. For closed-source models, we evaluate several OpenAI systems, including GPT-4o (Hurst et al., 2024), o3 mini, and o4 mini (OpenAI, 2025), and GPT-4.1 (OpenAI, 2025). We choose these models for their strong performance on a variety of code and reasoning benchmarks, and we evaluate whether code-specific training leads to measurable improvements. We also compare different model sizes, ranging from 3B and 4B to 32B parameters, in the open-source group
4.4 Evaluating Transfer Performance
To evaluate the extent to which the MetaLint training framework encourages adaptive reasoning over idiom specifications, as opposed to memorization, we construct a synthetic test set of 5.3k instances spanning 50 Ruff idioms, using the same data generation procedure described in Section 3.1. The data has a similar distribution of no-violation to violation examples as the SFT training set, with a 74:26 split in favor of no violations. We carefully select idioms to vary in their degree of overlap with those used for training, as shown in Figure 3. The selected idioms fall into three categories:
In-Domain: We select 5 idioms that are identical to those seen during SFT training, to assess whether MetaLint improves performance on idioms it was explicitly trained on.
Near Transfer: We include 10 idioms with specifications that closely resemble those in the training set but differ slightly. This category is crucial for probing memorization: if MetaLint relies too heavily on memorized training idioms, it may misclassify or underperform on these similar-looking test idioms due to interference.
Far Transfer: We choose 35 idioms that are substantially different from the training idioms to assess whether the LLM can rely solely on the provided specification. Success in this category would indicate that MetaLint fosters generalizable reasoning skills rather than dependence on specific training instances.
4.5 Evaluating Easy-to-Hard Generalization
To evaluate easy-to-hard generalization, we use the PEP hard idiom benchmark and compare the MetaLint-trained model against the baselines described in Section 4.3. The PEP idioms are held out for all models, including MetaLint, which is trained only on synthetic data generated from Ruff idioms. If training on the easier Ruff idioms leads to improved performance of MetaLint over the base models on the harder PEP idioms, it provides evidence that MetaLint encourages adaptive reasoning rather than memorization, thereby supporting easy-to-hard generalization.
5 Results
To test whether MetaLint enables the base LLM to interpret idiom specifications rather than simply memorize training examples, we evaluate its ability to generalize beyond the 50 Ruff idioms seen during training (Section 4.4) to both similarnear-transfer and significantly different far-transfer idioms. We also hypothesize that the RS-DPO (preference optimization) stage improves generalization over SFT alone, and that chain-of-thought (CoT) variants outperform their standard counterparts. Additionally, we assess easy-to-hard generalization by training on synthetic data for simpler Ruff idioms and evaluating on harder PEP idioms; if the model has truly learned to interpret high-level specifications, it should improve on these harder cases as well. Once again, we expect RS-DPO and CoT variants to yield stronger generalization. We present and analyze the results of these experiments in the following sections.
5.1 Ruff Idioms Performance
Table 1 presents Qwen3-4B’s idiom detection and localization performance on the full Ruff idiom transfer test set (Section 4.4) after MetaLint training, comparing CoT and non-CoT variants across both SFT and DPO stages. The untrained Qwen3-4B without CoT achieves decent detection precision but performs poorly on detection recall and localization. After SFT training, it shows substantial improvements in precision and all localization metrics, with modest recall gains. RS-DPO training further boosts detection recall and localization significantly (over 3× and nearly 5×, respectively), although with a slight drop in detection precision.
This precision drop stems from contrastive training examples where input $x$ has no violations, reducing the model’s accuracy in identifying such cases. We observe a strong correlation between the fraction of no-violation samples in training and detection precision. Including more such samples can improve precision but at the cost of recall and localization, leading to an unfavorable tradeoff. We find that including just 0.05% of no-violation data yields the best balance, minimizing precision loss while maximizing recall and improving detection F-score. Detailed ablation results are reported in Appendix A.2.2.
For the CoT variant, the untrained model already achieves strong detection precision and recall, along with high localization metrics—outperforming the SFT-only non-CoT model and trailing only slightly behind the SFT+RS-DPO non-CoT model (except in precision). To train the CoT model, we first apply RS-SFT to promote correct CoT reasoning and response formatting. This step is critical, as direct RS-DPO training alone causes metric collapse due to lack of absolute correctness supervision (since RS-DPO optimizes relative preference). RS-SFT leads to similar gains as non-CoT SFT, improving precision, recall, and localization. Final RS-SFT+RS-DPO training with 5% no-violation data mirrors the trends seen in the non-CoT case: increased recall, slight drop in precision, and overall gains in detection F-score and localization. Overall, the CoT model (RS-SFT+RS-DPO) is the strongest, achieving the highest precision and localization performance. However, the non-CoT SFT+RS-DPO model closely matches it in recall and approaches it on localization. Given that CoT inference is nearly 6× slower111measured on vLLM with 1× A100 80GB GPU the non-CoT model may be preferable in settings where detection recall and localization are the primary concerns.
5.2 Ruff Idioms Transfer Learning Performance
Table 2 presents MetaLint’s idiom detection performance across training setups for each transfer condition (in-domain, near-transfer and far-transfer) on the Ruff idiom benchmark. Untrained models perform poorly on in-domain idioms, especially in recall, suggesting these idioms are more subtle and difficult. Their stronger baseline performance on near and far transfer tasks likely reflects simpler or more surface-level idioms in those settings. SFT improves both precision and recall in-domain, but only precision improves in transfer settings, indicating limited generalization. Adding RS-DPO leads to consistent gains across all conditions, with particularly strong improvements in near and far transfer. This suggests that preference tuning encourages models to prioritize patterns that generalize beyond static rules.
CoT Untrained models already show high precision, likely due to the structured reasoning helping with correct interpretation of the idiom specifications and the output formatting instructions. RS-SFT improves recall with minimal changes to precision, though a small recall drop in near transfer may be due to a high baseline level of performance. Adding RS-DPO further boosts recall in all settings while maintaining strong precision. The largest recall improvements are seen in the in-domain setting, where idioms are hardest. These results offer some support for our claims that instruction-tuned models trained on synthetic linter signals generalize to novel idioms (the near and far transfer settings).
5.3 PEP Hard Idiom Benchmark Performance
Table 3 presents the results of evaluating the MetaLint DPO models and all baselines on the PEP hard idiom benchmark (a more comprehensive set of results is shown in Table 7). The MetaLint Qwen3-4B models significantly outperform the base Qwen3-4B models across all metrics in both the CoT and non-CoT settings, for both idiom violation detection and localization.
Interestingly, the non-CoT model achieves significantly higher detection recall and slightly higher F-score than the CoT variant, despite being less precise. Our analysis reveals that the CoT model’s lower detection recall results from its more conservative interpretation of the idiom specification and errors such as incorrect idiom interpretation, overthinking, and skipping lines, as detailed in Appendix A.3.3. In terms of detection F-score, the MetaLint models perform competitively, with only the closed-source GPT models, Qwen3-32B with CoT, and DeepSeek-R1-Distill-Qwen-32B with CoT surpassing the non-CoT model. Notably, the non-CoT model achieves the highest detection recall among all models evaluated, while the CoT model ranks among the best in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
For localization, the MetaLint models trail behind larger 32B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix A.3.2). This is notable given that the MetaLint models are much smaller (4B parameters) and trained only on synthetic data derived from easy idioms.
Overall, the strong performance, especially the best-in-class recall of the non-CoT model, demonstrates the effectiveness of our framework in achieving easy-to-hard generalization. This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction tuning, rather than relying on rote memorization.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Untrained 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
SFT+RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
CoT Untrained 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
CoT (RS-SFT) 0.935 0.8183 0.8727 0.6639 0.65 0.6569
CoT (RS-SFT + RS-DPO) 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 1: MetaLint Qwen3-4B results across training conditions evaluated on the Ruff meta-linting task.
Model In-Domain Near Transfer Far Transfer
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Untrained 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
SFT+RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
CoT Untrained 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
CoT (RS-SFT) 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
CoT (RS-SFT + RS-DPO) 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
Table 2: MetaLint Qwen3-4B detection results across training conditions evaluated on the Ruff meta-linting task across transfer conditions (In-Domain, Near Transfer and Far Transfer)
Detection Localization
Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
Qwen3-4B MetaLint (SFT+RS-DPO) (Ours) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) (Ours) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
Table 3: Results on the hard PEP benchmark to measure easy to hard generalization.
6 Conclusion
Our results demonstrate that MetaLint may encourage more adaptive reasoning over idiom specifications instead of rote memorization of code patterns through instruction-following and easy-to-hard generalization from synthetic data. By training on synthetic examples of easy-to-detect idioms derived from linter signals, MetaLint enables both Qwen3-4B CoT and non-CoT variants to generalize well across Ruff idioms in in-domain, near-transfer, and far-transfer settings, as well as to real-world, hard-to-detect PEP idioms. The models show substantial improvements in both detection and localization, particularly after RS-DPO training, with the non-CoT variant achieving the highest detection recall among all evaluated models, despite its smaller size and training solely on synthetic data. Additionally, it achieves this with just 4B parameters and no test-time compute in the non-CoT setting, making it one of the most efficient models. These findings underscore the effectiveness of synthetic instruction-tuned training for fostering generalization and reasoning, even in low-resource scenarios where annotated examples of harder idiom categories are limited or unavailable. Finally, we note that for easy idioms that are more “mechanical” to detect, linters remain the most cost-effective option. However, LLMs—when trained with the MetaLint approach—enable detection of more abstract idioms, potentially unlocking support for personalized linting, and offering a no-code solution for code quality enforcement.
7 Limitations and Future Work
Despite the promising results achieved by MetaLint, our work has several limitations that we plan to address in future research. First, our experiments are restricted to a single model family (Qwen3) and model size (4B). We intend to extend our evaluations to other model families, including LLaMA and Phi, as well as larger Qwen models (8B and 14B). In the CoT setting, it remains an open question whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model. We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data. As a result, we adopted a simpler rejection sampling strategy. Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as direct GRPO (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data. Our current experiments also focus exclusively on the Python programming language. Future work will explore extensions to other languages such as Java using Java-specific linters, as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa. Finally, while we prompt and train the models to generate refactoring fixes for code with idiom violations to improve code quality, we do not evaluate this capability in the present work and leave it for future exploration.
Author Contributions
Atharva Naik: Conception, Ruff Idiom Data Generation, SFT, RS-SFT, RS-DPO Experiments, Baselines, Writing
Lawanya Baghel: RS-DPO Experiments, PEP Hard Idiom Benchmark Creation, Writing
Dhakshin Govindarajan: PEP Hard Idiom Benchmark Creation
Darsh Agrawal: PEP Hard Idiom Benchmark Creation
Daniel Fried: Advising, Conception, Writing
Carolyn Rose: Advising, Conception, Writing
Acknowledgments
This work was partially supported by a generous collaborative research grant from Oracle Labs. We are grateful to Jack Sullivan, Qinlan Shen, Adam Pocock, and Nidhi Chandra at Oracle Labs for their guidance, feedback, and insightful discussions that played a key role in shaping the core ideas behind MetaLint and advancing the project.
Appendix A Appendix
A.1 Method Additional Details
A.1.1 MetaLint Instruction Following Prompt
We used the following instruction following style prompt to train the model with synthetic Ruff idiom data for the meta-linting task:
MetaLint Instruction Following Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} Violations per idiom:
An example input with the code file and idiom spec populated as well as the expected JSON style output is shown below:
Example Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations. Rationale: Type annotations are a good way to document the return types of functions. They also help catch bugs, when used alongside a type checker, by ensuring that the types of any returned values, and the types expected by callers, match expectation. Example:
def _add(a, b):
    return a + b
Use instead:
def _add(a: int, b: int) -> int:
    return a + b
Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say ’NO VIOLATIONS FOUND’ if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file:
  1 # -*- coding: utf-8 -*-
  2 # pragma pylint: disable=unused-argument, no-self-use
 86     def _reload(self, event, opts):
 87         """Configuration options have changed,
    save new values"""
 88         self.options = opts.get("fn_cisco_amp4ep", {})
 89         validate_opts(self)
 90
 91     @function("fn_amp_move_computer")
 92     def _fn_amp_move_computer_function(self, event, *args,
    **kwargs):
 93         """Function: Move computer to a group with given
    connector guid and group guid."""
 94         try:
Violations per idiom:
Example Output **Idiom ANN202 Violations:**
{"line": " 86     def _reload(self, event, opts):", "fix": null}
{"line": " 92     def _fn_amp_move_computer_function(self,
event, *args, **kwargs):", "fix": null}
A.1.2 DPO Contrastive Pair and RS-SFT Sampling Details
To generate RS-DPO contrastive samples (or RS-SFT outputs) from the baseline SFT (or untrained) models, we used the following hyperparameters: nucleus sampling with a maximum of 2048 new tokens, $k=5$ sampled outputs per input, temperatures picked cyclically from {0, 0.3, 0.5, 0.7, 1}, a top-$p$ (cumulative probability threshold) of 0.95, and a seed of $42+i$, where $i\in\{1,\dots,k\}$, to encourage both reproducibility and output diversity.
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded “Baseline Inference Prompt” described in Section A.1.4.
A.1.3 Training Hyperparameters and Computational Environment
We use the following hyperparameters for the SFT and DPO training experiments.
SFT/RS-SFT hyperparameters:
For both SFT and RS-SFT we fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision. The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled to reduce memory usage, with non-reentrant mode. Evaluation is performed every 2000 steps, and checkpoints are saved at the same interval. Special tokens are manually handled in the chat template without automatic insertion. The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
RS-DPO parameters:
We fine-tune the model using RS-DPO with bfloat16 precision and a reward shaping parameter $\beta=0.1$. Training is performed for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and a warmup ratio of 0.1. We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled with non-reentrant mode to optimize memory usage. The optimizer is AdamW, and evaluation is conducted every 200 steps with checkpoints saved at the same interval. The training is seeded with 42 for reproducibility.
Computational Environment: All SFT, RS-SFT and RS-DPO experiments were conducted on a Linux server equipped with two NVIDIA A100 80GB PCIe GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03. Each job had access to 100 GB of CPU memory and 2 CPU cores. Training was performed using mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage. For inference we used a similar environment but varied number of GPUs based on model sizes and needs.
A.1.4 Baseline Inference Details
We use the following hyperparameters for performing inference with the baseline LLMs:
Open Source LLMs: We perform nucleus sampling with 8192 max-new tokens, temperature of 0.7, top-p (cumulative probability threshold) of 0.95 and seed of 42 (to promote reproducibility).
Closed Source LLMs: We use the chat completion OpenAI API with max tokens of 1024 for GPT-4.1 and GPT-4o and max completion tokens of 3000 for o3-mini and o4-mini. We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty).
Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches. For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Idiom Benchmark.
Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called “### Final Idiom Violations Found”. Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
[…]
Inquiry.objects.filter(owner=kwargs.pop(’user’)))", "fix": null}
[…]
4 from inquiry.models import *",
[…]
and as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:** NO VIOLATIONS FOUND Violations per idiom:
A.2 Additional Experimental Details
A.2.1 Idioms Chosen for Ruff Idiom Transfer Dataset
Table 4 lists the Ruff idioms used in the SFT training and synthetic transfer evaluation test sets. Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
Training Set Idioms Test Set Idioms
PyFlakes: PyFlakes:
F405, F501, F502, F601, F621 F403, F406, F503, F602, F622
pycodestyle: pycodestyle:
E402, E701, E721, E741, E743 E401, E702, E722, E731, E742
Naming: Miscellaneous:
N801, N802, N803, N804, N805, ERA001, C901, I001, I002, BLE001
N806, N807, N811, N812, N813 (shared with training)
pyupgrade: flake8 annotations:
UP001, UP002, UP003, UP004, UP005, UP006, ANN001, ANN002, ANN003, ANN201, ANN202,
UP007, UP008, UP009, UP010, UP011, ANN204, ANN205, ANN206
UP040, UP044, UP045, UP046, UP047
Miscellaneous: flake8 async:
ERA001, C901, I001, I002, BLE001 ASYNC100, ASYNC105, ASYNC109, ASYNC110,
ASYNC115, ASYNC116, ASYNC210, ASYNC220,
ASYNC221, ASYNC222, ASYNC230, ASYNC251
Bugbear: flake8 bandit:
B002, B003, B004, B005, B006, S102, S103, S104, S105, S106,
B007, B008, B009, B010, B012 S107, S108, S110, S112, S113,
S201, S202, S301, S302, S303
Table 4: Ruff idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
Figure 3: ID: In-Domain, NeT: Near Transfer, FaT: Far Transfer.
A.2.2 DPO No Violation Fraction Ablations
We analyze the impact of varying the amount of samples with zero violations used for RS-DPO training. These experiments were motivated by initial findings comparing models trained only on data with at least one violation to those trained on the full dataset. By design, RS-DPO generates significantly more training data for cases with at least one violation, due to greater variance in reward signals. This is further amplified by the fact that the initial SFT policy/checkpoint is already quite accurate in handling cases with NO VIOLATIONS FOUND leading to low variance in reward across responses.
Our early experiments showed that excluding all NO VIOLATIONS FOUND cases led to notable gains in recall and line-level localization. However, this came at the cost of a significant drop in precision compared to the SFT policy/base model. Further analysis revealed a sharp decline in the accuracy of predicting NO VIOLATIONS FOUND, from nearly 99% down to 70-80%, with performance worsening monotonically over training steps. Conversely, training on the full dataset (i.e., including 100% of the NO VIOLATIONS FOUND cases) improved precision but offered only modest gains in recall and localization, which also degraded with continued training. These findings suggest that while some NO VIOLATIONS FOUND data is necessary to maintain high precision, too much of it may hinder recall and localization.
To investigate this trade-off, we experimented with keeping only a fraction of the NO VIOLATIONS FOUND data during training. Specifically, we randomly sampled $k$of such data, varying $k$ across {0%, 2%, 5%, 10%, 20%, 40%, 100%}. These percentages were selected based on observed trends: 20%, 40%, and 100% yielded similar results, which discouraged further tests at 60% or 80%, while 2% and 5% were chosen due to a noticeable performance jump between 0% and 10%. We found that 5% offered a favorable middle ground, largely retaining or slightly reducing precision, while preserving most of the recall (resulting in the highest detection F-score), and only modestly impacting line-level localization. Based on these insights, we conducted a limited ablation on the CoT model, evaluating 2% and 5% inclusion to determine the optimal setting for both detection and localization (as shown in Table 6).
Fraction of NV data Det. Line overlap (Det. + Loc.)
P R F P R F
0% 0.6268 0.9577 0.7577 0.6777 0.6932 0.6854
2% 0.671 0.9128 0.7734 0.6681 0.6812 0.6746
5% 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
10% 0.7584 0.8114 0.784 0.6263 0.6474 0.6367
20% 0.8382 0.7227 0.7762 0.5721 0.5815 0.5768
40% 0.8683 0.5618 0.6822 0.4683 0.4735 0.4709
100% 0.8565 0.4152 0.5593 0.4041 0.4056 0.4048
Table 5: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT. Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND. Conversely, including 100% improves precision but leads to reduced recall and localization performance. All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Idiom Transfer test set.
Fraction of NV data Det. Line overlap (Det. + Loc.)
P R F P R F
2% 0.9226 0.8901 0.906 0.7688 0.7638 0.7663
5% 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 6: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model with CoT. We perform limited ablations because of the insights from the non CoT model training.
A.3 More Results
A.3.1 Expanded Results on the PEP Hard Idiom Benchmark
We show the expanded results across various model sizes for the evaluated model families in Table 7. We note that most results follow the expected trends with more parameters or CoT usage leading to better performance but there are soem exceptions to the trend. We mainly see this for cases like Qwen2.5 and Qwen2.5Coder families. We note that Qwen2.5Coder-7B-Instruct has almost zero metrics because it always predicts NO VIOLATIONS FOUND for all instances and Qwen2.5Coder-14B-Instruct has really low scores because of similar reasons. for Qwen2.5 family we notice that 32B variant performs a bit worse than 32B.
We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models. This suggests that SFT alone may lead to overfitting on the Ruff idiom distribution and struggles to generalize from easy to hard cases without DPO training. These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline. However, we also emphasize that while the SFT stage can limit generalization, it remains essential for effective DPO training, as it teaches the LLM to follow the correct output format and establishes a strong base policy. This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
Detection Localization
Model P R F P R F
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-3B-Instruct 0.0667 0.0033 0.0063 0.0036 0.0036 0.0036
Qwen2.5-7B-Instruct 0.4333 0.1379 0.2092 0.0585 0.0518 0.0549
Qwen2.5-14B-Instruct 0.8017 0.4324 0.5618 0.2389 0.2158 0.2267
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
Qwen2.5Coder-3B-Instruct 0.7802 0.411 0.5384 0.1257 0.0745 0.0936
Qwen2.5Coder-7B-Instruct 0.0667 0.0033 0.0063 0 0 0
Qwen2.5Coder-14B-Instruct 0.2 0.0443 0.0726 0.0294 0.0264 0.0278
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-7B with CoT 0.7143 0.2841 0.4065 0.1064 0.1122 0.1092
DeepSeek-R1-Distill-Qwen-14B with CoT 0.69 0.2345 0.35 0.1856 0.1245 0.149
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
MetaLint (SFT) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
MetaLint w CoT (RS-SFT) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
Table 7: Result on the hard PEP benchmark to measure easy to hard generalization.
A.3.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
To analyze the statistical significance of performance differences over the PEP benchmark, we conduct Wilcoxon signed-rank tests comparing various MetaLint variants against each other and against baseline models. We evaluate instance-level detection accuracy (binary labels indicating whether the LLM correctly predicted the presence of a violation) as well as instance-level precision and recall for line-level localization. To control for multiple comparisons, we apply a Bonferroni correction to adjust the significance threshold $\alpha$ as $\alpha=\frac{0.05}{m}$ where $m$ is the number of comparisons (or rows in any given statistical significance table in this case).
Table 9 reports the Wilcoxon signed-rank test statistic and corresponding $p$-value (in parentheses) for detection accuracy, localization precision, and localization recall when comparing various MetaLint variants to assess the effects of RS-DPO and CoT. We find that applying RS-DPO to the base SFT policy leads to statistically significant improvements in both detection and localization performance, with RS-DPO consistently outperforming the original SFT checkpoint across all three metrics with it being always better for localization. For the CoT variant, RS-DPO also yields consistent but less significant gains, likely because the RS-SFT CoT checkpoint is already relatively strong. Finally, we observe no statistically significant difference between the CoT (RS-SFT+RS-DPO) and the standard (SFT+RS-DPO) variant, suggesting that CoT does not provide a meaningful additional benefit in this setting.
Table 10 shows the statistical significance of comparing the base untrained model Qwen3-4B with its MetaLint variants (SFT and SFT+RS-DPO), and the Qwen3-4B CoT model with MetaLint w/ CoT (RS-SFT and RS-SFT+RS-DPO). The SFT variant yields significant gains in detection and localization recall, but not in localization precision. The SFT+RS-DPO model improves significantly across all three metrics. In contrast, training RS-SFT from the Qwen3-4B w/ CoT base does not yield significant improvements. However, the RS-SFT+RS-DPO variant produces significant gains in localization precision and recall, but not detection. These results suggest that while SFT alone offers limited generalization, combining it with DPO reliably improves localization and can significantly boost detection when starting from a weaker base model.
Table 11 shows the statistical significance results when comparing the MetaLint (SFT+RS-DPO) and MetaLint w CoT (RS-SFT+RS-DPO) variants against various baselines. Here we want to higlight that MetaLint offers comparable performance across two out of three or all three metrics against several 32B models that outperform it like Qwen3-32B, Qwen3-32B w CoT, Qwen2.5Coder-32B and R1-Distill-Qwen-32B. Also the MetaLint non CoT (SFT+RS-DPO) variant has no singificant difference in performance compared to o3-mini, soldifying that MetaLint without CoT has generalized to the point of being as capable as o3-mini (even though the Qwen3-4B models without CoT and Qwen3-4B model with CoT perform worse than it with the difference being statistically singificant in Table 8).
Model Comparison Detection Localization P Localization R
Qwen3-4B vs o3-mini 1266.5 (7.20e-21) 743.5 (2.96e-11) 739.0 (4.23e-09)
Qwen3-4B w CoT vs o3-mini 921.5 (3.23e-09) 2385.5 (9.61e-02) 1891.0 (4.99e-04)
Table 8: Wilcoxon signed-rank test results comparing untrained Qwen3-4B variants with o3-mini, using Bonferroni-adjusted significance threshold $\alpha=0.025$. Each cell reports the test statistic (p-value).
Table 12 shows the effect of using a CoT for the Qwen3 model families and we notice that using a CoT leads to singificant gains for all metrics for the 4B and 8B models indicating that for smaller models CoTs might be essential for good performance on this task. However the 14B and 32B model only show statistically significant improvement in localization precision with the CoT indicating that the CoT might offer limited benefit for larger models.
Table 13 shows the effect of varying model scale for the Qwen3, Qwen2.5, Qwen2.5Coder, and DeepSeek-R1-Distill-Qwen families. For Qwen3 we see benefits moving from 4B to 8B abd 8B to 14B but no statistically significant difference moving from 14B to 32B when not using a CoT. Wehn using a CoT for Qwen3 we notice that the performance differences are rarely different in terms of statistical significant except for localizaiton performance between 4B and 8B and 8B and 14B. For R1-Distill-Qwen family we notice a significant difference moving from 14B to 32B but not for 7B to 14B. For the Qwen2.5Coder family we notice difference across all model scales, but the trend is weird with a big drop in performance from 3B to 7B and then a slow climb back to great performance around 32B. We notice that for the Qwen2.5 family which shows relatively reasonable trends with model scale, the performance differences are statistically singificant execpt for the performance gain from 14B to 32B being significant only for recall. To conclude the trends across model scales vary a lot across model families but in general the model size does help but differences may be smaller if the models are capable of reasoning and use a CoT.
Table 14 shows comparison between the GPT models. We only compared GPT-4o and its successor GPT-4.1 and o3-mini against o4-mini and the results show that GPT-4.1 is only significantly better for localization recall while o4-mini is beter than o3-mini for overall localization but not for detection.
Model Comparison Detection Localization P Localization R
MetaLint (SFT) vs MetaLint (SFT+RS-DPO) 7192.0 (1.92e-12) 0.0 (2.49e-20) 0.0 (2.92e-18)
MetaLint w CoT (RS-SFT) vs MetaLint w CoT (RS-SFT+RS-DPO) 839.5 (2.18e-03) 740.0 (3.95e-03) 523.0 (2.34e-05)
MetaLint (SFT) vs MetaLint w CoT (RS-SFT) 528.0 (6.91e-14) 11.0 (1.38e-15) 113.0 (7.95e-12)
MetaLint (SFT+RS-DPO) vs MetaLint w CoT (RS-SFT+RS-DPO) 8140.0 (5.55e-01) 2568.5 (8.42e-01) 2544.0 (4.44e-01)
Table 9: Wilcoxon signed-rank test results comparing MetaLint variants. Each cell reports test statistic (p-value). All the MetaLint models are trained Qwen3-4B variants. We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B MetaLint (SFT) 560.5 (8.64e-03) 363.5 (1.82e-02) 238.0 (4.85e-04)
Qwen3-4B vs Qwen3-4B MetaLint (SFT+RS-DPO) 7260.0 (4.13e-09) 411.0 (1.99e-15) 979.5 (7.30e-09)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT) 1224.0 (7.22e-01) 937.0 (6.12e-01) 918.5 (5.39e-01)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) 1728.0 (1.83e-02) 1011.0 (1.53e-03) 966.0 (1.02e-03)
Table 10: Wilcoxon signed-rank test results comparing MetaLint models against their untrained counterparts, with Bonferroni-adjusted significance threshold $\alpha=0.0125$. Each cell reports the test statistic (p-value).
Model Comparison Detection Localization P Localization R
Qwen3-8B vs MetaLint (SFT+RS-DPO) 8140.5 (7.21e-03) 1309.5 (2.19e-08) 2067.0 (2.12e-03)
Qwen3-8B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2070.0 (9.17e-01) 1974.5 (1.88e-01) 2161.0 (6.58e-01)
Qwen3-14B vs MetaLint (SFT+RS-DPO) 7304.0 (4.96e-01) 2816.5 (3.15e-02) 3159.0 (2.88e-02)
Qwen3-14B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2392.0 (2.78e-01) 2749.5 (1.26e-01) 2319.0 (1.32e-03)
Qwen3-32B vs MetaLint (SFT+RS-DPO) 7175.0 (4.48e-01) 3262.0 (1.93e-02) 2818.5 (5.56e-03)
Qwen3-32B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2677.5 (9.95e-03) 3479.0 (8.38e-02) 3180.5 (5.64e-04)
R1-Distill-Qwen-7B vs MetaLint (SFT+RS-DPO) 8244.0 (2.65e-08) 555.0 (8.12e-15) 1924.0 (1.91e-04)
R1-Distill-Qwen-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 2907.0 (7.07e-10) 915.5 (1.04e-12) 1569.5 (8.64e-06)
R1-Distill-Qwen-14B vs MetaLint (SFT+RS-DPO) 9877.0 (3.99e-06) 2582.5 (9.36e-06) 3085.0 (2.00e-03)
R1-Distill-Qwen-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 2660.0 (9.11e-08) 1703.0 (2.98e-06) 1791.0 (3.97e-05)
R1-Distill-Qwen-32B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 5767.5 (1.93e-01) 3705.0 (6.67e-06)
R1-Distill-Qwen-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 3125.0 (3.11e-02) 3641.5 (6.35e-02) 2175.0 (4.99e-06)
Qwen2.5-3B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (1.24e-22) 68.5 (1.26e-19)
Qwen2.5-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (4.24e-22) 0.0 (9.89e-20)
Qwen2.5-7B vs MetaLint (SFT+RS-DPO) 7312.5 (3.37e-10) 208.0 (1.44e-18) 610.0 (1.99e-12)
Qwen2.5-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 1187.5 (1.14e-14) 226.0 (2.60e-18) 406.5 (5.69e-14)
Qwen2.5-14B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 3045.5 (5.70e-04) 4006.0 (3.83e-01)
Qwen2.5-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 4123.0 (4.86e-01) 3228.0 (1.51e-03) 4383.0 (9.89e-01)
Qwen2.5-32B vs MetaLint (SFT+RS-DPO) 8640.0 (1.76e-04) 1492.5 (3.12e-08) 2971.5 (4.55e-02)
Qwen2.5-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1792.0 (8.16e-06) 1166.0 (2.43e-08) 1983.5 (4.71e-03)
Qwen2.5Coder-3B vs MetaLint (SFT+RS-DPO) 11184.0 (8.64e-03) 953.5 (3.73e-12) 1716.0 (3.00e-07)
Qwen2.5Coder-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 3683.5 (6.45e-03) 1126.0 (4.39e-11) 1403.5 (1.32e-08)
Qwen2.5Coder-7B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (7.69e-23) 0.0 (2.02e-20)
Qwen2.5Coder-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (2.61e-22) 0.0 (6.55e-20)
Qwen2.5Coder-14B vs MetaLint (SFT+RS-DPO) 9123.5 (1.04e-12) 289.0 (8.19e-20) 736.0 (7.32e-14)
Qwen2.5Coder-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 1112.0 (1.82e-19) 159.5 (7.63e-20) 408.5 (3.02e-15)
Qwen2.5Coder-32B vs MetaLint (SFT+RS-DPO) 6833.5 (2.86e-01) 4500.0 (9.59e-01) 2651.5 (7.07e-05)
Qwen2.5Coder-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1039.5 (1.16e-02) 2655.0 (9.39e-01) 1235.5 (1.44e-05)
o3-mini vs MetaLint (SFT+RS-DPO) 7520.0 (4.83e-02) 4986.0 (5.20e-01) 4427.5 (2.87e-01)
o3-mini vs MetaLint w CoT (RS-SFT+RS-DPO) 1944.0 (7.15e-04) 3169.0 (5.16e-01) 2683.0 (4.23e-01)
Table 11: Wilcoxon signed-rank test statistics and p-values comparing MetaLint variants against baseline models. All the MetaLint variants are Qwen3-4B variants and Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0017$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B w CoT 1260.0 (3.99e-08) 425.0 (1.91e-09) 624.5 (2.25e-05)
Qwen3-8B vs Qwen3-8B w CoT 1924.0 (4.27e-03) 1132.5 (8.69e-06) 1005.0 (1.15e-04)
Qwen3-14B vs Qwen3-14B w CoT 1691.0 (2.01e-01) 2127.0 (4.27e-04) 2398.5 (1.26e-01)
Qwen3-32B vs Qwen3-32B w CoT 1572.5 (2.75e-01) 1767.0 (6.35e-06) 2596.5 (7.31e-02)
Table 12: Wilcoxon signed-rank test results measuring the effect of Chain-of-Thought (CoT) prompting across Qwen3 model scales. Each cell reports test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-8B 546.0 (2.35e-08) 618.0 (1.37e-04) 572.5 (1.72e-03)
Qwen3-8B vs Qwen3-14B 1350.0 (2.11e-03) 1503.5 (3.96e-05) 1008.5 (2.69e-07)
Qwen3-14B vs Qwen3-32B 875.0 (2.22e-02) 3129.5 (7.94e-01) 2061.0 (4.14e-01)
Qwen3-4B w CoT vs Qwen3-8B w CoT 1468.5 (1.90e-02) 1578.5 (1.07e-01) 1081.5 (2.60e-03)
Qwen3-8B w CoT vs Qwen3-14B w CoT 904.5 (1.40e-01) 1248.0 (1.66e-03) 1099.5 (1.82e-05)
Qwen3-14B w CoT vs Qwen3-32B w CoT 850.0 (3.78e-02) 1834.5 (4.87e-01) 2396.0 (4.40e-01)
R1-Distill-Qwen-7B vs R1-Distill-Qwen-14B 4278.0 (2.67e-01) 1431.0 (4.66e-03) 1962.5 (6.01e-01)
R1-Distill-Qwen-14B vs R1-Distill-Qwen-32B 2432.0 (1.44e-12) 1475.5 (8.07e-11) 843.5 (6.38e-15)
Qwen2.5Coder-3B vs Qwen2.5Coder-7B 1541.0 (4.56e-14) 0.0 (3.46e-10) 0.0 (7.07e-10)
Qwen2.5Coder-7B vs Qwen2.5Coder-14B 8.0 (7.89e-04) 0.0 (2.04e-03) 0.0 (2.14e-03)
Qwen2.5Coder-14B vs Qwen2.5Coder-32B 423.0 (2.83e-27) 100.5 (2.00e-22) 43.5 (9.29e-23)
Qwen2.5-3B vs Qwen2.5-7B 18.0 (2.43e-08) 0.0 (3.43e-05) 0.0 (3.58e-05)
Qwen2.5-7B vs Qwen2.5-14B 960.5 (1.70e-13) 500.5 (1.65e-11) 574.0 (2.62e-10)
Qwen2.5-14B vs Qwen2.5-32B 1925.0 (1.87e-04) 1858.5 (6.60e-03) 2108.5 (1.43e-02)
Table 13: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings. Each cell shows the test statistic (p-value). All Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
Model Comparison Detection Localization P Localization R
GPT-4o vs GPT-4.1 2550.0 (9.21e-01) 3961.0 (5.83e-01) 2207.5 (4.48e-06)
o3-mini vs o4-mini 575.0 (6.68e-01) 1171.0 (1.44e-03) 1079.0 (1.47e-04)
Table 14: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
A.3.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
We observe that a significant portion of the lower detection recall of the CoT MetaLint Qwen3-4B model, relative to its non CoT counterpart, can be attributed to its higher tendency to predict NO VIOLATIONS FOUND in cases that do, in fact, contain violations. Specifically, the CoT model fails to flag violations in 89 additional instances compared to the non CoT model, amounting to nearly 17% of the evaluation set (89 out of 536 examples).
The idiom wise distribution of these missed violations is shown in Figure 4. While the failure distribution follows a somewhat long tail pattern, the most significant drops occur for PEP 614, PEP 616, and PEP 593. Notably, if the CoT model matched the non CoT model’s performance on just these three PEPs, its detection recall would rise to 0.605, surpassing that of all open source baselines evaluated.
Upon inspecting CoT traces for these and other idioms (see examples in Table 15), we identify several recurring failure modes: 1) Ambiguity in interpreting the idiom specification. For example, in PEP 614, which targets decorators with complex expressions, the CoT model often labels expressions that humans consider complex as simple. 2) Overthinking and repetitive reasoning traces, particularly for PEP 616. 3) Skipping or entirely missing lines that contain violations, again observed in PEP 616. 4) Underspecified idioms. For instance, in PEP 593, which recommends using the Annotated type from the typing module to attach metadata to type hints, the spec lacks clarity and concrete examples, making it hard to learn what constitutes a violation.
We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__. The CoT model often misclassifies such “simple” use cases as complex.
Overall, these patterns suggest that the CoT model applies the idiom specifications more conservatively, resulting in higher precision but at the cost of reduced recall.
Figure 4: Distribution of comparative failures of the CoT MetaLint Qwen3-4B model relative to its non-CoT variant. While errors span a long tail across many PEPs, the majority are concentrated in three: PEP614, PEP593, and PEP616, which motivates our focused analysis on these cases.
PEP Chain of Thought Reason for Failure
614
… I need to check the code file for violations of the 614 idiom,
which is about using complex expressions in decorators. …
… The register variable is defined on line 7 as partial(register,
server=hippa). So, the decorator is a partial function. …
… the idiom allows for complex expressions in the decorator.
For example, if the decorator was something like
@lambda func: timed(…)(func), that would be allowed. …
… the register is a partial function, which is a simple expression. …
… the decorator is just a variable that’s a partial function. …
Ambiguity about what constitutes
a complex expression. The idiom
targets decorators with complex
expressions that make code
more verbose
525
… The user … wants me to check for … idiom, 525, which
is about using asynchronous generators with ”async for”
syntax … the violation would be if the code uses
synchronous generators (like regular for loops) instead of
asynchronous generators (using async for) … Wait, the
example given in the idiom is a class that uses __aiter__
and __anext__ to create an asynchronous iterator. The fix
is to use a function with async def and yield. So, I need to
check if there are any instances where the code is using
synchronous generators (like for loops) instead of
asynchronous ones. … Looking through the code, I
don’t see any for loops that are using ’async for’. …
CoT accidentally looks for the
incorrect construct at a crucial
point (async for instead of for)
487
… using __init_subclass__ and __set_name__ for class
customization without metaclasses … using metaclasses
for simple customization tasks makes code unnecessarily
complex … there’s a class called FirestoreClientMeta,
which is a metaclass … used for registration of transports
… the problem is whether this is a simple customization …
not just for a single customization but for managing
multiple transports … the idiom is about using metaclasses
for simple customization … but in this case … handling a
more complex task … maybe this is not a violation …
Ambiguity around what constitutes
a simple class. The idiom targets
instances where metaclasses
are used for simple customization
tasks that could have been handled
with ”__init_subclass__” or
”__set_name__”)
Table 15: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.

### v2

@@
 Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices.
-We introduce MetaLint, a new instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications.
-Unlike conventional approaches that train models on static, rule-based data, MetaLint employs instruction tuning on synthetic linter-generated data to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining.
+We introduce MetaLint, an instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications.
+Unlike conventional approaches that train models on static code quality conventions, MetaLint employs instruction tuning on synthetic linter-generated data with dynamic conventions to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining.
 To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize.
-Our results show that MetaLint improves generalization to unseen PEP idioms, achieving a 70.37% F-score on idiom detection with the highest recall (70.43%) among all evaluated models.
-It also achieves 26.73% on localization, competitive for its 4B parameter size and comparable to larger state-of-the-art models like o3-mini, highlighting its potential for future-proof code quality analysis.
+Our results show that MetaLint training improves generalization to unseen idioms.
+Qwen3-4B attains a 70.37% F-score on a manually curated and challenging PEP idiom detection benchmark, achieving the highest recall (70.43%) among all evaluated models.
+For localization, it reaches 26.73%, which is a strong outcome for its 4B parameter size and comparable to larger state-of-the-art models such as o3-mini, highlighting its potential for future-proof code quality analysis.
+Furthermore, MetaLint training enables generalization in idiom detection across model families, model scales, synthetic data from diverse linters, and Java idioms, demonstrating the general applicability of our approach.
+We plan to release our code and data to enable reproducibility and further work.
 1 Introduction
-With the rise of Large Language Model (LLM) agents, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent.
-Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023).
-However, many of these efforts are hindered by reliance on task-specific training data, often based on narrow or outdated coding best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024).
+With the rise of Large Language Models (LLM) of code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent Singhal et al.
+(2024); Zheng et al.
+(2024).
+Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Singhal et al., 2024; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023).
+Several evaluation studies indicate that LLMs struggle with this task Singhal et al.
+(2024); Zheng et al.
+(2024), while attempts to improve them through prompting or training are limited by task-specific, static datasets often grounded in narrow or outdated coding practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b).
 As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024).
 They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024).
-Ideally, we would develop LLM systems capable of identifying code quality issues without explicit supervision for the target idioms, especially for hard or rare patterns, and that can adapt to evolving best practices over time.
-One approach to this problem is to train the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms.
+Ideally, we would develop LLM systems that can identify code quality issues without explicit supervision for target idioms—especially hard or rare patterns—and adapt to evolving best practices over time.
+We approach this problem by training the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms.
 For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function.
-However, as noted in PEP 506 (D’Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against the usage of this module for security reasons, which is often missed by developers as highlighted by accepted answers on forums like StackOverflow.
+However, as noted in PEP 506 (D'Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against using this module for security reasons, which is often missed by developers, as highlighted by accepted answers on forums like StackOverflow.
 PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom.
 As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024).
-Additionally, the issues can be fixed by replacing instances of “bad” idioms with corresponding “good” idioms that align with best practices.
-In this work we extended this observation by training the LLM on a higher-level instruction-following task: given a specification of a best-practice code idiom $I$, the model learns to identify and localize code fragments that violate the idiom and could be rewritten to adhere to or incorporate idiom $I$.
-We refer to this instruction-following task as meta-linting.
-The core idea is to teach the LLM the underlying structure of the task, namely identifying and locating idioms based on high-level specifications, rather than training it to memorize or learn a static set of best practices.
-Since the focus is on improving the model’s instruction-following ability rather than tailoring it to a particular data distribution of best practices, this approach can keep up with evolving best practices.
-Our approach, MetaLint, is motivated by the observation that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022).
-Since meta linting frames each idiom as a separate task, instruction style training is naturally suited to generalize to novel and unseen idioms.
-Existing linters for programming languages (e.g., Ruff for Python) implement numerous simple rules, or “easy” idioms, that reflect aspects of code quality such as modern syntax, readability, and security.
-These linters can be used to generate large-scale synthetic data for supervised instruction fine-tuning and serve as verifiers during preference optimization or reinforcement learning.
-This setup also supports easy-to-hard generalization: training on automatically generated linter data for easy idioms helps build a strong model that can handle harder, more abstract idioms at test time, which go beyond the reach of static analyzers.
-Similar strategies have shown success in math reasoning tasks (Sun et al., 2024b).
-To systematically study this generalization, we construct a challenging benchmark of hard idioms based on popular PEPs that introduce high-level constructs.
-We evaluate state-of-the-art reasoning and code models on this benchmark and compare them to models trained with MetaLint, focusing on whether LLMs can move beyond memorization, generalize from easy to hard idioms, and adapt to novel patterns across code quality domains.
-Our key contributions are as follows:
+Additionally, these issues can be addressed by replacing instances of ``bad'' idioms with corresponding ``good'' idioms that align with best practices.
+Moreover, for this example and similar abstract idioms, constructing a precise rule-based approach is difficult.
+Simply flagging any use of random.choice, even in non-security-critical scenarios (e.g., randomization in a game engine), could result in a poor user experience.
+Vijayvergiya et al.
+(2024) show that LLMs can capture abstract notions of code quality, such as code idioms where building a linter or rule-based approach is challenging, by incorporating semantic reasoning about code and developer intent.
+In this work, we train LLMs to recognize code idioms through a higher-level instruction-following task dubbed ``meta-linting'': given a specification of a best-practice code idiom $I$, the model learns to identify and localize non-idiomatic code fragments.
+Our pipeline is designed to support easy-to-hard generalization (Sun et al., 2024b).
+The easy cases involve simple idioms that can already be captured by existing linters, while the hard cases correspond to nuanced patterns such as PEP 506, where constructing precise rule-based checks is infeasible.
+To enable this, we generate synthetic training data for easy idioms using available linters and leverage it to improve performance on harder cases where linter support is lacking.
+While prior work such as Zhang et al.
+(2024c; b) has explored automated refactoring of non-idiomatic Python code, including the use of LLMs with prompting, our focus differs in three ways.
+First, we target challenging idioms beyond the reach of current linters.
+Second, we train on easy idioms with the goal of transferring detection ability to harder cases.
+Finally, we emphasize adaptability, aiming for LLMs that can accommodate evolving best practices provided in-context as instructions and examples, rather than memorizing a static rule sets.
+To tackle meta-linting, we introduce MetaLint, a training framework motivated by prior work showing that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022).
+Since meta-linting treats each idiom as a distinct task or code quality judgment, instruction fine-tuning (IFT) and preference optimization (PO) naturally extend detection ability to novel idioms.
+Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing simple idioms, which we use both for supervised IFT and as verifiers during PO to improve performance on harder idioms.
+To systematically study this generalization, we construct a benchmark of challenging idioms derived from popular PEPs introducing high-level constructs.
+We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy idioms.
+Our key contributions are:
   1. 1.
-    We introduce MetaLint, an instruction-following framework that leverages synthetic data for easy-to-hard generalization and addresses the limitations of static training used in prior LLM-based code quality approaches.
+    We introduce MetaLint, a training framework that leverages instruction following and synthetic data to enable easy-to-hard generalization while remaining adaptable to evolving best practices.
   2. 2.
-    We construct a benchmark of challenging, widely relevant idioms inspired by various PEPs, and evaluate state-of-the-art code and reasoning models on it.
+    We construct a benchmark of challenging, broadly relevant code-quality idioms inspired by PEPs to evaluate the extent of easy-to-hard generalization achieved by MetaLint.
   3. 3.
-    We analyze whether MetaLint promotes adaptive reasoning over memorization and enables generalization from easy to hard idioms.
+    We benchmark state-of-the-art code and reasoning models on our PEP hard-idiom benchmark and compare them against MetaLint-trained models.
+Our method achieves the highest detection recall and competitive localization scores, even with smaller 4B models and without test-time compute.
+  4. 4.
+    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), test-time reasoning settings (with and without CoT), and model scales (3B–8B).
 2 Related Work
-2.1 Code Quality Analysis with Large Language Models
+Code Quality Analysis with Large Language Models.
 A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis.
-Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as linters guided by rule descriptions, relying on prompting or fine-tuning to improve rule enforcement.
-LiteLLM (Fang et al., 2025) leverages LLMs to reduce both the cost and false positive rate of linting for Verilog.
+Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as rule-guided linters via prompting or fine-tuning.
+While Blyth et al.
+(2025) proposes a static analysis-driven prompting framework to improve LLM-generated code, Du et al.
+(2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives.
+LintLLM (Fang et al., 2025) and (Shin et al., 2025) leverages LLMs for linting of Verilog and Quantum computing code.
 Khare et al.
-(2023) show that LLMs can outperform traditional static analysis tools for detecting certain security-related CWEs, especially when step-by-step reasoning is used.
+(2023) show LLMs outperform traditional static analysis tools for security-related CWEs with step-by-step reasoning.
 Vijayvergiya et al.
-(2024) train LLMs to detect and localize violations of best practices, while Rasheed et al.
-Rasheed et al.
-(2024) design a multi-agent review pipeline targeting issues like maintainability, efficiency, and bugs.
-Other works (Jiang et al., 2025b; Yao et al., 2025) use comparative prefix-tuning and reinforcement learning with static analysis–based rewards to promote the generation of higher-quality functionally correct code.
+(2024) train LLMs for best practice violation detection and localization, while Rasheed et al.
+(2024) design a multi-agent review pipeline for maintainability, efficiency, and bugs.
+Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation.
 Naik et al.
@@
 (2025) integrate LLMs with linters to produce more informative code reviews.
-Finally, CoUpJava (Jiang et al., 2025a) presents a benchmark of Java version–related code upgrade histories, conceptually similar to the hard PEP idiom benchmark we developed for Python.
-While such efforts highlight the potential of LLMs in code quality tasks, they often target specific rule sets or best practices, necessitating retraining as those evolve.
-In contrast, our instruction-following meta-linting framework trains models to interpret high-level code quality specifications and perform static analysis to enforce them, enabling broader generalization beyond fixed rules.
-2.2 Instruction Following for Generalization
-Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning a single fixed task.
-Prior work shows that exposing models to diverse task instructions allows them to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022).
+RIdiom (Zhang et al., 2024c) introduces a rule-based way to identify and refactor non-idiomatic Python code with AST rewrite rules, while Zhang et al.
+(2024b) combines LLMs and rule-based detectors but doesn't explore nuanced idioms like PEP 506 or training LLMs to keep up with evolving best practices.
+Finally, CoUpJava (Jiang et al., 2025a) presents Java version upgrade benchmarks, conceptually similar to our hard PEP idiom benchmark for Python.
+Although prior work demonstrates the potential of LLMs for code quality tasks, it focuses on fixed rule sets or best practices that require retraining as they evolve.
+In contrast, we train models to interpret high-level specifications and perform static analysis, enabling broader generalization.
+Instruction Following for Generalization.
+Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning fixed tasks.
+Prior work shows diverse task instructions allow models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022).
 Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020).
-Instructions serve as high-density task representations, acting as a substitute for supervision (Puri et al., 2022) and enabling models to generalize even with minimal labeled data or via pseudo-labeled examples (Gu et al., 2022).
-Studies also show that instruction diversity is a key driver of generalization, with varied instructions improving performance more than repeated exposure to the same task format (Charton et al., 2024).
-This phenomenon holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023).
-SELF-GUIDE (Zhao et al., 2024) also performs task-specific instruction following using synthetic data, demonstrating its effectiveness; however, it relies entirely on LLM-generated data and does not incorporate verifiers.
-Collectively, these results suggest that instruction tuning functions as task-level meta-learning, equipping models with the ability to rapidly adapt to new tasks via natural language.
-Inspired by this work we treat modeling of specific code quality idioms as individual tasks and propose the MetaLint training framework that generates synthetic data at scale for each meta-task to facilitate cross-idiom generalization.
-This way the trained model can keep up with new idioms and evolving best practices.
-2.3 Easy-to-Hard Generalization
-Recent research demonstrates that easy-to-hard generalization improves model performance across mathematical, algorithmic, and coding tasks, motivating its application to code quality analysis.
-In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025).
-Several works implement curriculum strategies to structure this progression (Parashar et al., 2025; Zhao et al., 2025), while others emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024).
-These principles extend beyond math: Sun et al.
-(2024a) shows that reward models trained on simple code and math problems improve performance on complex ones, and Zhao et al.
-(2025) further explores generalization in code tasks through curriculum fine-tuning.
-Multi-task and length generalization studies (Hu et al., 2025) and differentiable programming setups (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code.
-Zhang et al.
-(2024) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats.
-While our approach does not explicitly follow curriculum learning, we draw inspiration from this line of work by training MetaLint on large-scale synthetic data that captures easily detectable code idioms, many of which are directly covered by rule-based linters.
-We hypothesize that these simple patterns act as a stepping stone, enabling generalization to more complex and abstract unseen PEP idioms.
+Instructions serve as high-density task representations, substituting supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022).
+Studies also show that instruction diversity drives generalization, with varied instructions outperforming repeated exposure to identical formats (Charton et al., 2024).
+This phenomenon holds across domains, including program synthesis where task-level prompting facilitates generalization in code generation models (Niu et al., 2023).
+SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness, but relying entirely on LLM-generated data without verifiers.
+These results suggest instruction tuning acts as task-level meta-learning, enabling models to adapt to new tasks through natural language.
+Building on this we model specific code quality idioms as individual tasks and generate large-scale synthetic data for each meta-task to support cross-idiom generalization.
+This allows the trained model to keep pace with new idioms and evolving best practices.
+We also discuss additional related work on easy-to-hard generalization in Appendix B.
 3 Method
-Figure 1: MetaLint training framework (1/2): Core components of our framework include using linter tools (like Ruff) for 1) Synthetic Data Generation, 2) Supervised Instruction Fine-Tuning (SFT) with synthetic data and a 3) Verifiable Reward Model (also derived from the Ruff linter).
-Figure 2: MetaLint training framework (2/2): We use the linter-based verifiable reward model for the final two preference optimization components: 4) Rejection Sampling Direct Preference Optimization (RS-DPO), and, 5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
-We propose the MetaLint framework to support the learning of skills essential for code idiom analysis, such as interpreting idiom descriptions in terms of underlying code constructs and performing line-by-line static analysis to detect idiom violations.
-We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}={D_{I},E_{I}}$.
+We design the MetaLint framework to teach an LLM to operationalize idiom descriptions provided in context, rather than memorizing specific idioms, thereby enabling adaptation to novel idioms at test time.
+We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$.
 The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$.
@@
 3.1 Synthetic Data Generation
-A key advantage of our meta-task formulation is its support for easy-to-hard generalization.
-We train LLMs on a set of idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$ to improve their ability to flag idiom violations, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by $\mathcal{L}$).
-Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand, and detect code idioms from high level descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to an untrained checkpoint.
+One of the main goals of our meta-task formulation is enabling easy-to-hard generalization.
+We train LLMs on a set of ``easy'' idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter).
+Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand and detect code idioms from in-context descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to the untrained model.
 Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision.
-For this purpose, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, semantics, and more.
-We run Ruff on Python source code files $f\in\mathcal{F}$ from the STACK-V2 (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories.
+For Python, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, etc., while for Java, we use the PMD static analyzer, which covers 269 idioms as well as some manually written tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/ queries inspired by 8 Java Enhancement Protocols (JEPs) (Table 8).
+We run Ruff, PMD, and the JEP tree-sitter queries on Python and Java source code files $f\in\mathcal{F}$ from the STACK (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories.
 This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$.
 Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data.
-Additionally, the Ruff website hosts rule-specific documentation, including descriptions and examples, which we scrape to construct meta-task prompts $M_{I_{\mathcal{L}}}$ for each idiom.
-An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix A.1.1.
+Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples.
+An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix C.1.
+For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.
 3.2 Instruction Supervised Fine-Tuning
-As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of 50 linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$.
+As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$.
 The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$.
-The model’s output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix A.1.1).
+The model's output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix C.1).
 In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND.
-We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some idioms, the final distribution is approximately 70:30 in favor of files with no violations.
-This results in a total of 53k training instances spanning 50 idioms.
+We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python data, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data.
+This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
 3.3 Verifiable Reward Model and Preference Optimization
@@
 We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
-Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter’s line numbers (yellow circle in “Verifiable Reward Model”, Figure 1) as ground truth and the model’s predicted lines (blue circle) as predictions.
+Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter's line numbers (blue circle in ``Verifiable Reward Model'', Figure 1) as ground truth and the model’s predicted lines (yellow circle) as predictions.
 Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap.
@@
 For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$.
-Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
+Following Khaki et al.
+(2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
 $\Phi^{RL}=\arg\max\sum_{(x,y_{win},y_{lose})\in\mathcal{D}_{p}}\log\sigma\left(\beta\log\frac{\Phi^{RL}(y_{win}|x)}{\Phi^{SFT}(y_{win}|x)}-\beta\log\frac{\Phi^{RL}(y_{lose}|x)}{\Phi^{SFT}(y_{lose}|x)}\right)$
@@
 Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning.
-To obtain CoT traces that help the LLM arrive at the correct answer, we adopt a rejection sampling approach (“Rejection Sampling SFT” in Figure 2) for collecting SFT data.
-For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained LLM capable of CoT reasoning (e.g., Qwen3-4B), and compute the reward $r_{y_{i}}$ for each response, following a similar procedure as in RS-DPO.
-However, instead of forming contrastive pairs, we discard any $y_{i}$ with reward $r_{y_{i}}<\gamma$, where $\gamma=1$.
-In other words, we discard CoT–response pairs that are incorrect or improperly formatted and retain only those that are fully correct.
-The reward is applied only to the final response, obtained after extracting or parsing the CoT trace.
-We also remove any cases where the CoT fails to terminate or produce a final answer.
-For some inputs $x$, no valid $y_{i}$ may be found.
-In such cases, we skip the input entirely.
-To promote meta-task diversity, we retain at most two valid responses $y_{i}$ per input.
-We sample multiple $y_{i}$ only for inputs with at least one violation, and only a single $y_{i}$ for inputs with no violations, this helps maintain a similar no-violation-to-violation ratio of 71:29 as in the regular SFT data.
-The latter case is also more likely to yield a valid CoT–response pair within the token limit.
-When more valid responses are available than needed, we keep the shortest $y_{i}$ completions, as they typically reflect more concise reasoning (the final answer usually consumes a similar number of tokens across samples).
-Following this policy, we collect 52.7k instances, which we use to train the base model using SFT.
-After training the LLM on this dataset, we obtain a CoT-capable SFT model $\Phi_{CoT}^{SFT}$.
-We then apply the same RS-DPO procedure described in Section 3.3, with the only modification being that each $y_{i}$ now contains both the CoT trace and the final response.
+To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection.
+For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained CoT-capable LLM (e.g., Qwen3-4B), and compute a reward $r_{y_{i}}$ for each, following the RS-DPO procedure in Figure 2.
+Instead of forming contrastive pairs, we discard any $y_{i}$ with $r_{y_{i}}<\gamma$, where $\gamma=1$, i.e., CoT–response pairs that are incorrect or improperly formatted.
+Rewards are applied only to the final response, obtained after parsing the CoT trace, and we also remove cases where the CoT fails to terminate or yield an answer.
+If no valid $y_{i}$ is found for an input $x$, we skip it.
+To promote meta-task diversity, we retain at most two valid responses per input: multiple $y_{i}$ only for violation cases and a single $y_{i}$ otherwise.
+This maintains the 71:29 no-violation-to-violation ratio of Ruff Python SFT data, with the latter more likely to fit within token limits.
+When excess valid responses exist, we keep the shortest completions, as they typically reflect more concise reasoning (final answers are of similar token length across samples).
+Following this policy, we collect 52.7k Python training instances from Ruff data, which we use to train the reasoning-enabled base Qwen3-4B with SFT.
+This yields a CoT-capable SFT model $\Phi_{CoT}^{SFT}$ for Python code quality analysis.
+We then apply the RS-DPO procedure in Section 3.3 and Figure 2, with the only change being that each $y_{i}$ now includes both the CoT trace and final response.
+Figure 1: MetaLint: (1) Synthetic data generation with linters/tools, (2) Supervised Instruction Fine-Tuning (SFT) on this data, and (3) Verifiable Reward Model derived from the linter.
+Figure 2: MetaLint: Preference Optimization using reward model: (4) Rejection Sampling Direct Preference Optimization (RS-DPO), and (5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
 4 Experiments
-4.1 PEP Hard Idiom Benchmark
-To evaluate whether MetaLint training helps the base LLM learn to interpret high-level idiom specifications, and apply this skill to harder, more abstract idioms that linters struggle to detect, we construct a benchmark of “hard idioms” sourced from PEPs.
-We begin by identifying fifteen PEPs that define semantic or abstract behaviors beyond what syntax or token patterns capture, then design specific heuristics for each to detect code that violates their guidelines.
-Using these heuristics, we search the STACK‑V2 corpus for violations, prioritizing high recall over strict precision to retrieve a broad set of candidates for manual selection.
-These “hard idioms” cannot be reliably detected by simple pattern matching, making them ideal for testing whether models truly understand idiom specifications or memorize frequent patterns seen in the Ruff idioms.
-From the retrieved examples, we handpick 15–20 representative files per PEP that clearly illustrate each violation.
-For each file, we identify the exact lines showing the violation (“before” snippet) and author a corrected version (“after” snippet) that complies with the PEP, forming the ground truth.
-To balance the benchmark, we include negative examples (code without violations) by selecting files retrieved for one PEP and confirming they don’t violate others.
-This results in 536 challenging examples, with 52% containing at least one violation and 48% being violation-free, allowing us to evaluate MetaLint’s ability to generalize from easy to hard idioms.
 4.1 Evaluation Metrics
 We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs.
-For both tasks, we compute precision, recall, and F-score metrics.
-Detection metrics are calculated at the corpus level for each idiom, treating each idiom as a distinct class.
-Since the distribution of idiom occurrences may be imbalanced, we use macro averaging across idioms to weigh each class equally.
-Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations.
-For each instance $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$, we define the indicator variable:
-$\mathds{1}[x]=\begin{cases}1&\text{if }x\text{ is true}\\ 0&\text{otherwise}\end{cases}$
-The idiom detection metrics per idiom $I$ are defined as:
-$P_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|=0]\cdot\mathds{1}[|\hat{y}|>0]\right)}$
-$R_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|=0]\right)}$
-We compute the final detection metrics by macro averaging across all idioms:
-$P_{\text{Det}}=\frac{1}{|I|}\sum_{I}P_{I},\quad R_{\text{Det}}=\frac{1}{|I|}\sum_{I}R_{I},\quad F_{\text{Det}}=\frac{2P_{\text{Det}}R_{\text{Det}}}{P_{\text{Det}}+R_{\text{Det}}}$
-For idiom localization, we compute set level precision, recall, and F1-score for each instance $(x,y)\in\mathcal{D}$, comparing the predicted and ground truth violation lines:
-$P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
-4.3 Baselines
-We evaluate a range of state-of-the-art open and closed-source LLMs, trained for coding and reasoning tasks, on the PEP hard idiom benchmark to contextualize the improvements achieved by the MetaLint framework in flagging idiom violations from high-level specifications.
-Among open source models, we include instruction-tuned variants from the Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), and Qwen3 (Team, 2025) families.
-For closed-source models, we evaluate several OpenAI systems, including GPT-4o (Hurst et al., 2024), o3 mini, and o4 mini (OpenAI, 2025), and GPT-4.1 (OpenAI, 2025).
-We choose these models for their strong performance on a variety of code and reasoning benchmarks, and we evaluate whether code-specific training leads to measurable improvements.
-We also compare different model sizes, ranging from 3B and 4B to 32B parameters, in the open-source group
-4.4 Evaluating Transfer Performance
-To evaluate the extent to which the MetaLint training framework encourages adaptive reasoning over idiom specifications, as opposed to memorization, we construct a synthetic test set of 5.3k instances spanning 50 Ruff idioms, using the same data generation procedure described in Section 3.1.
-The data has a similar distribution of no-violation to violation examples as the SFT training set, with a 74:26 split in favor of no violations.
-We carefully select idioms to vary in their degree of overlap with those used for training, as shown in Figure 3.
-The selected idioms fall into three categories:
-In-Domain: We select 5 idioms that are identical to those seen during SFT training, to assess whether MetaLint improves performance on idioms it was explicitly trained on.
-Near Transfer: We include 10 idioms with specifications that closely resemble those in the training set but differ slightly.
-This category is crucial for probing memorization: if MetaLint relies too heavily on memorized training idioms, it may misclassify or underperform on these similar-looking test idioms due to interference.
-Far Transfer: We choose 35 idioms that are substantially different from the training idioms to assess whether the LLM can rely solely on the provided specification.
-Success in this category would indicate that MetaLint fosters generalizable reasoning skills rather than dependence on specific training instances.
-4.5 Evaluating Easy-to-Hard Generalization
-To evaluate easy-to-hard generalization, we use the PEP hard idiom benchmark and compare the MetaLint-trained model against the baselines described in Section 4.3.
-The PEP idioms are held out for all models, including MetaLint, which is trained only on synthetic data generated from Ruff idioms.
-If training on the easier Ruff idioms leads to improved performance of MetaLint over the base models on the harder PEP idioms, it provides evidence that MetaLint encourages adaptive reasoning rather than memorization, thereby supporting easy-to-hard generalization.
+For both tasks, we report precision, recall, and F-score metrics.
+Detection metrics are calculated at the corpus level for each idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers.
+To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics).
+For localization, metrics are averaged only across instances with at least one line of idiom violation in the ground truth.
+Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix D.1.
+4.2 Generalization on Synthetic Data
+To evaluate whether MetaLint training produces adaptive LLMs that handle evolving best practices and novel idioms at test time, we explore transfer settings spanning Python & Java.
+Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1.
+The data has a 74:26 no-violation-to-violation split, similar to the SFT training set.
+Idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
+In domain. 5 idioms identical to those in SFT training, to assess whether MetaLint improves performance on explicitly trained idioms.
+Near transfer. 10 idioms with specifications similar but not identical to training idioms to probe memorization. Reliance on memorized patterns, may hurt performance due to interference.
+Far transfer. 35 idioms distinct from training, to test whether the LLM can follow the provided specification and adapt to novel idioms at test time.
+For these experiments, we use Qwen3-4B (with and without reasoning) and Llama-3.2-3B-Instruct to study the effect of test-time compute and model family.
+PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries.
+We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer between PMD and JEP idioms to test adaptation to novel Java idioms.
+These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
+4.3 PEP Hard Idiom Benchmark
+Benchmark Construction: To test whether MetaLint helps LLMs interpret high-level idiom specifications and generalize to nuanced idioms that linters miss, we construct a benchmark of ``hard idioms'' from 15 PEPs defining semantic or abstract behaviors beyond syntax.
+We design heuristics per PEP (Table 13, 14 and 15) to detect guideline violations and search the STACK-V2 corpus, prioritizing recall to retrieve broad candidate sets for manual selection.
+These idioms cannot be reliably detected by simple pattern matching, making them ideal for evaluating model's true understanding versus rote memorization.
+From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (“start” and “end”) of the violated code, providing ground-truth for localization.
+We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases.
+The final benchmark contains 536 examples (52% violations, 48% violation-free), enabling evaluation of MetaLint’s generalization from easy to hard Python idioms.
+Evaluating Easy-to-Hard Generalization: We use the PEP hard idiom benchmark to test whether training on synthetic data for linter-detectable idioms improves performance on hard idioms.
+We evaluate the base model, SFT, and DPO-trained models on this benchmark.
+Benchmarking on Hard Idioms: We evaluate state-of-the-art open and closed-source code and reasoning LLMs on the PEP hard idiom benchmark, comparing them to MetaLint-trained models.
+Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B Agarwal et al.
+(2025).
+Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 OpenAI (2025b).
+We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
 5 Results
-To test whether MetaLint enables the base LLM to interpret idiom specifications rather than simply memorize training examples, we evaluate its ability to generalize beyond the 50 Ruff idioms seen during training (Section 4.4) to both similarnear-transfer and significantly different far-transfer idioms.
-We also hypothesize that the RS-DPO (preference optimization) stage improves generalization over SFT alone, and that chain-of-thought (CoT) variants outperform their standard counterparts.
-Additionally, we assess easy-to-hard generalization by training on synthetic data for simpler Ruff idioms and evaluating on harder PEP idioms; if the model has truly learned to interpret high-level specifications, it should improve on these harder cases as well.
-Once again, we expect RS-DPO and CoT variants to yield stronger generalization.
-We present and analyze the results of these experiments in the following sections.
-5.1 Ruff Idioms Performance
-Table 1 presents Qwen3-4B’s idiom detection and localization performance on the full Ruff idiom transfer test set (Section 4.4) after MetaLint training, comparing CoT and non-CoT variants across both SFT and DPO stages.
-The untrained Qwen3-4B without CoT achieves decent detection precision but performs poorly on detection recall and localization.
-After SFT training, it shows substantial improvements in precision and all localization metrics, with modest recall gains.
-RS-DPO training further boosts detection recall and localization significantly (over 3× and nearly 5×, respectively), although with a slight drop in detection precision.
-This precision drop stems from contrastive training examples where input $x$ has no violations, reducing the model’s accuracy in identifying such cases.
-We observe a strong correlation between the fraction of no-violation samples in training and detection precision.
-Including more such samples can improve precision but at the cost of recall and localization, leading to an unfavorable tradeoff.
-We find that including just 0.05% of no-violation data yields the best balance, minimizing precision loss while maximizing recall and improving detection F-score.
-Detailed ablation results are reported in Appendix A.2.2.
-For the CoT variant, the untrained model already achieves strong detection precision and recall, along with high localization metrics—outperforming the SFT-only non-CoT model and trailing only slightly behind the SFT+RS-DPO non-CoT model (except in precision).
-To train the CoT model, we first apply RS-SFT to promote correct CoT reasoning and response formatting.
-This step is critical, as direct RS-DPO training alone causes metric collapse due to lack of absolute correctness supervision (since RS-DPO optimizes relative preference).
-RS-SFT leads to similar gains as non-CoT SFT, improving precision, recall, and localization.
-Final RS-SFT+RS-DPO training with 5% no-violation data mirrors the trends seen in the non-CoT case: increased recall, slight drop in precision, and overall gains in detection F-score and localization.
-Overall, the CoT model (RS-SFT+RS-DPO) is the strongest, achieving the highest precision and localization performance.
-However, the non-CoT SFT+RS-DPO model closely matches it in recall and approaches it on localization.
-Given that CoT inference is nearly 6× slower111measured on vLLM with 1× A100 80GB GPU the non-CoT model may be preferable in settings where detection recall and localization are the primary concerns.
-5.2 Ruff Idioms Transfer Learning Performance
-Table 2 presents MetaLint’s idiom detection performance across training setups for each transfer condition (in-domain, near-transfer and far-transfer) on the Ruff idiom benchmark.
-Untrained models perform poorly on in-domain idioms, especially in recall, suggesting these idioms are more subtle and difficult.
-Their stronger baseline performance on near and far transfer tasks likely reflects simpler or more surface-level idioms in those settings.
-SFT improves both precision and recall in-domain, but only precision improves in transfer settings, indicating limited generalization.
-Adding RS-DPO leads to consistent gains across all conditions, with particularly strong improvements in near and far transfer.
-This suggests that preference tuning encourages models to prioritize patterns that generalize beyond static rules.
-CoT Untrained models already show high precision, likely due to the structured reasoning helping with correct interpretation of the idiom specifications and the output formatting instructions.
-RS-SFT improves recall with minimal changes to precision, though a small recall drop in near transfer may be due to a high baseline level of performance.
-Adding RS-DPO further boosts recall in all settings while maintaining strong precision.
-The largest recall improvements are seen in the in-domain setting, where idioms are hardest.
-These results offer some support for our claims that instruction-tuned models trained on synthetic linter signals generalize to novel idioms (the near and far transfer settings).
-5.3 PEP Hard Idiom Benchmark Performance
-Table 3 presents the results of evaluating the MetaLint DPO models and all baselines on the PEP hard idiom benchmark (a more comprehensive set of results is shown in Table 7).
-The MetaLint Qwen3-4B models significantly outperform the base Qwen3-4B models across all metrics in both the CoT and non-CoT settings, for both idiom violation detection and localization.
-Interestingly, the non-CoT model achieves significantly higher detection recall and slightly higher F-score than the CoT variant, despite being less precise.
-Our analysis reveals that the CoT model’s lower detection recall results from its more conservative interpretation of the idiom specification and errors such as incorrect idiom interpretation, overthinking, and skipping lines, as detailed in Appendix A.3.3.
-In terms of detection F-score, the MetaLint models perform competitively, with only the closed-source GPT models, Qwen3-32B with CoT, and DeepSeek-R1-Distill-Qwen-32B with CoT surpassing the non-CoT model.
-Notably, the non-CoT model achieves the highest detection recall among all models evaluated, while the CoT model ranks among the best in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
-For localization, the MetaLint models trail behind larger 32B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix A.3.2).
-This is notable given that the MetaLint models are much smaller (4B parameters) and trained only on synthetic data derived from easy idioms.
-Overall, the strong performance, especially the best-in-class recall of the non-CoT model, demonstrates the effectiveness of our framework in achieving easy-to-hard generalization.
-This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction tuning, rather than relying on rote memorization.
+To test whether MetaLint training leads to cross-idiom generalization instead of mere memorization of the training idioms and whether it can produce models that can keep up with evolving code quality standards, we present the transfer performance on the synthetic data for ``easy'' idioms in section 5.1.
+Then we explore the extent to which MetaLint training achieves easy-to-hard generalization from the synthetic easy idioms to hard, manually curated PEP idioms in section 5.2.
+Finally, we compare MetaLint trained models against state-of-the-art code and reasoning models on the manually curated hard PEP idioms in section 5.3.
+5.1 Generalization on Synthetic Data
+Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 18).
+While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 16.
+The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop in the case of Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics at the cost of a slight drop in detection precision.
+We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix D.3.
+Additionally, Table 16 shows that while SFT can lead to slight gains for the transfer settings (near transfer and far transfer), most gains emerge in the DPO stage, especially for non-reasoning models and detection recall.
+Overall this suggests that SFT can lead to memorization of the training idioms while DPO leads to generalization to novel idioms.
 Model Detection Localization
 $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
-Untrained 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
-SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
-SFT+RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
-CoT Untrained 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
-CoT (RS-SFT) 0.935 0.8183 0.8727 0.6639 0.65 0.6569
-CoT (RS-SFT + RS-DPO) 0.9234 0.8643 0.8929 0.771 0.7571 0.764
-Table 1: MetaLint Qwen3-4B results across training conditions evaluated on the Ruff meta-linting task.
-Model In-Domain Near Transfer Far Transfer
-$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
-Untrained 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
-SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
-SFT+RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
-CoT Untrained 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
-CoT (RS-SFT) 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
-CoT (RS-SFT + RS-DPO) 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
-Table 2: MetaLint Qwen3-4B detection results across training conditions evaluated on the Ruff meta-linting task across transfer conditions (In-Domain, Near Transfer and Far Transfer)
+Qwen3-4B 0.5380 0.2637 0.3539 0.1396 0.1479 0.1436
+Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.2960 0.2968
+Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
+Qwen3-4B w CoT 0.8812 0.6854 0.7710 0.5049 0.4878 0.4962
+Qwen3-4B w CoT + RS-SFT 0.9350 0.8183 0.8727 0.6639 0.6500 0.6569
+Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.7710 0.7571 0.7640
+Table 1: Cross-Idiom Generalization on Python Ruff Idioms: Effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning).
+Best score across the compared training setups per model are bolded.
+PMD and JEP Tree-Sitter Idioms: To demonstrate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 28).
+Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO.
+For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct.
+Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP.
+In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores because our metrics exclude that class for detection and only score positive cases for localization.
+Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5).
+In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD’s broader diversity and more challenging specifications (Appendix C.5).
+Overall, MetaLint training consistently yields more adaptable models than the base model, but performance depends on the diversity of training idioms and the gap in instruction quality between training and test data.
+Model Transfer Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
+Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
+Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.2800
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.5930 0.5969 0.5949
+Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
+Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
+Table 2: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28).
+The transfer column indicates training and test data on the left and right side of the arrow.
+Best score across the compared training setups per model are bolded.
+5.2 Evaluating Easy-to-Hard Generalizations
+To evaluate whether MetaLint training on easy, linter-detectable Ruff idioms improves performance on hard, manually curated PEP idioms, we report results on our PEP hard idiom benchmark (Table 3, full results in Table 19).
+At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT can induce memorization of the training distribution and reduce adaptability.
+In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix E.2).
+An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited.
+Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to unlock easy-to-hard generalization.
+Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision.
+Our analysis attributes the CoT model’s reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix E.3.
+Model Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
+Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
+Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
+Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
+Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.4970 0.2785 0.1437 0.1896
+Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+Table 3: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3).
+Best score across the compared training setups per model are bolded.
+5.3 Benchmarking on Hard Idioms
+Table 4 compares the best-performing Qwen3-4B MetaLint DPO models against state-of-the-art code and reasoning models (full results in Table 17).
+Detection: In terms of detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini).
+However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
+Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix E.2) and outperform GPT-oss-20B.
+This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy idioms, and that the non-CoT model does not use test-time compute.
+Overall, the strong results, especially the best-in-class recall of the non-CoT model, demonstrate the effectiveness of our framework in achieving easy-to-hard generalization.
+This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction fine-tuning and DPO rather than relying on rote memorization.
 Detection Localization
 Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
-Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
-Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
 Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
 Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
-Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
+Qwen3-14B 0.9021 0.4612 0.6103 0.2890 0.2521 0.2693
 Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
 Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
-Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
-Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
+Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.3540
+Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.1630 0.1477 0.1550
 Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
-DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
-Qwen3-4B MetaLint (SFT+RS-DPO) (Ours) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
-Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) (Ours) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
-GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
-GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
+DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.7130 0.4015 0.3403 0.3684
+GPT-oss-20b 0.8377 0.3531 0.4968 0.2510 0.1695 0.2024
+GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
+Qwen3-4B MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
+Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
 o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
 o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
-Table 3: Results on the hard PEP benchmark to measure easy to hard generalization.
-6 Conclusion
-Our results demonstrate that MetaLint may encourage more adaptive reasoning over idiom specifications instead of rote memorization of code patterns through instruction-following and easy-to-hard generalization from synthetic data.
-By training on synthetic examples of easy-to-detect idioms derived from linter signals, MetaLint enables both Qwen3-4B CoT and non-CoT variants to generalize well across Ruff idioms in in-domain, near-transfer, and far-transfer settings, as well as to real-world, hard-to-detect PEP idioms.
-The models show substantial improvements in both detection and localization, particularly after RS-DPO training, with the non-CoT variant achieving the highest detection recall among all evaluated models, despite its smaller size and training solely on synthetic data.
-Additionally, it achieves this with just 4B parameters and no test-time compute in the non-CoT setting, making it one of the most efficient models.
-These findings underscore the effectiveness of synthetic instruction-tuned training for fostering generalization and reasoning, even in low-resource scenarios where annotated examples of harder idiom categories are limited or unavailable.
-Finally, we note that for easy idioms that are more “mechanical” to detect, linters remain the most cost-effective option.
-However, LLMs—when trained with the MetaLint approach—enable detection of more abstract idioms, potentially unlocking support for personalized linting, and offering a no-code solution for code quality enforcement.
-7 Limitations and Future Work
-Despite the promising results achieved by MetaLint, our work has several limitations that we plan to address in future research.
-First, our experiments are restricted to a single model family (Qwen3) and model size (4B).
-We intend to extend our evaluations to other model families, including LLaMA and Phi, as well as larger Qwen models (8B and 14B).
-In the CoT setting, it remains an open question whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model.
-We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data.
-As a result, we adopted a simpler rejection sampling strategy.
-Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as direct GRPO (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data.
-Our current experiments also focus exclusively on the Python programming language.
-Future work will explore extensions to other languages such as Java using Java-specific linters, as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa.
-Finally, while we prompt and train the models to generate refactoring fixes for code with idiom violations to improve code quality, we do not evaluate this capability in the present work and leave it for future exploration.
-Author Contributions
-Atharva Naik: Conception, Ruff Idiom Data Generation, SFT, RS-SFT, RS-DPO Experiments, Baselines, Writing
-Lawanya Baghel: RS-DPO Experiments, PEP Hard Idiom Benchmark Creation, Writing
-Dhakshin Govindarajan: PEP Hard Idiom Benchmark Creation
-Darsh Agrawal: PEP Hard Idiom Benchmark Creation
-Daniel Fried: Advising, Conception, Writing
-Carolyn Rose: Advising, Conception, Writing
+GPT-4o 0.8938 0.6788 0.7716 0.4461 0.3320 0.3807
+GPT-4.1 0.9070 0.6460 0.7546 0.4632 0.4673 0.4653
+GPT-5 (high) 0.9130 0.5673 0.6998 0.4397 0.4257 0.4326
+Table 4: Benchmarking on Hard Idioms: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training.
+The best scores are bolded and second best and underlined.
+6 Conclusion and Future Work
+Our results show that MetaLint training fosters adaptive reasoning over idiom specifications rather than rote memorization.
+We observe generalization to unseen idioms in Python and Java, across three linters (Ruff, PMD, JEP tree-sitter), two model families (Qwen, Llama), reasoning and non-reasoning settings, and multiple scales (3B, 4B, 8B).
+Easy-to-hard generalization occurs from linter-detectable Ruff idioms to harder PEP idioms, with SFT teaching output formatting and DPO enabling true generalization.
+Compared to state-of-the-art code and reasoning models, MetaLint-trained Qwen models have detection comparable with o3-mini and GPT-5, achieving highest recall (non-CoT) and third-best precision (CoT).
+Localization lags but surpasses GPT-oss-20B with only 4B parameters and no test-time compute and is comparable to o3-mini, demonstrating efficiency.
+These results highlight the effectiveness of instruction fine-tuning and preference optimization on synthetic data for reasoning and generalization, even with scarce annotated examples.
+For mechanically easy idioms, linters remain cost-effective, but MetaLint enables detection of abstract idioms, supporting personalized, evolving code quality standards.
+We plan to release code and data for reproducibility.
+Future work includes training for automated refactoring and exploring advanced RL methods like Group Relative Policy Optimization (GRPO) Shao et al.
+(2024).
 Acknowledgments
@@
 We are grateful to Jack Sullivan, Qinlan Shen, Adam Pocock, and Nidhi Chandra at Oracle Labs for their guidance, feedback, and insightful discussions that played a key role in shaping the core ideas behind MetaLint and advancing the project.
-Appendix A Appendix
-A.1 Method Additional Details
-A.1.1 MetaLint Instruction Following Prompt
+Appendix A Limitations
+Despite the promising results achieved by MetaLint, our work has some limitations that we plan to address in future research.
+For the CoT setting, we didn't explore whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model.
+We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data.
+As a result, we adopted a simpler rejection sampling or RS-SFT strategy.
+Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data.
+Our current experiments also focus on training on one language at a time, such as only Python or Java.
+Future work will explore joint training and extension to more programming languages like JavaScript, Ruby, Go, etc., as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa.
+Finally, while we do not evaluate or train for refactoring of the idiom-violating code, we plan to do so in future work.
+Appendix B More Related Work
+Easy-to-Hard Generalization.
+Research shows that training on simpler problems enhances generalization to harder ones in math, algorithms, and code, motivating its application to code quality analysis.
+In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025).
+Several works emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024).
+Beyond math, Sun et al.
+(2024a) shows that reward models trained on simple code and math problems improve performance on complex ones.
+Broader studies on multi-task and length generalization (Hu et al., 2025) and differentiable programming (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code.
+Zhang et al.
+(2024a) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats.
+Drawing inspiration from this work, we train MetaLint on large-scale synthetic data covering easily detectable code idioms handled by rule-based linters, and hypothesize that these simple patterns serve as stepping stones toward generalizing to complex, novel PEP idioms.
+Appendix C Method Additional Details
+C.1 MetaLint Instruction Following Prompt
 We used the following instruction following style prompt to train the model with synthetic Ruff idiom data for the meta-linting task:
@@
 An example input with the code file and idiom spec populated as well as the expected JSON style output is shown below:
-Example Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations.
+Example Ruff Meta-Task Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations.
 Rationale: Type annotations are a good way to document the return types of functions.
@@
 You should also suggest a fix if possible.
-Report the results per idiom specification mentioned above and just say ’NO VIOLATIONS FOUND’ if no violations are found for a given idiom.
+Report the results per idiom specification mentioned above and just say 'NO VIOLATIONS FOUND' if no violations are found for a given idiom.
 Do not detect any idioms not specified above.
@@
 Violations per idiom:
-Example Output **Idiom ANN202 Violations:**
+Example Ruff Meta-Task Output **Idiom ANN202 Violations:**
 {"line": " 86     def _reload(self, event, opts):", "fix": null}
@@
 event, *args, **kwargs):", "fix": null}
-A.1.2 DPO Contrastive Pair and RS-SFT Sampling Details
+C.2 DPO Contrastive Pair and RS-SFT Sampling Details
 To generate RS-DPO contrastive samples (or RS-SFT outputs) from the baseline SFT (or untrained) models, we used the following hyperparameters: nucleus sampling with a maximum of 2048 new tokens, $k=5$ sampled outputs per input, temperatures picked cyclically from {0, 0.3, 0.5, 0.7, 1}, a top-$p$ (cumulative probability threshold) of 0.95, and a seed of $42+i$, where $i\in\{1,\dots,k\}$, to encourage both reproducibility and output diversity.
 For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models.
-In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded “Baseline Inference Prompt” described in Section A.1.4.
-A.1.3 Training Hyperparameters and Computational Environment
-We use the following hyperparameters for the SFT and DPO training experiments.
-SFT/RS-SFT hyperparameters:
-For both SFT and RS-SFT we fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision.
+In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section C.4.
+C.3 Training Hyperparameters and Computational Environment
+Python SFT/RS-SFT hyperparameters:
+We fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision.
 The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1.
@@
 The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
-RS-DPO parameters:
+Python RS-DPO parameters:
 We fine-tune the model using RS-DPO with bfloat16 precision and a reward shaping parameter $\beta=0.1$.
@@
 The training is seeded with 42 for reproducibility.
-Computational Environment: All SFT, RS-SFT and RS-DPO experiments were conducted on a Linux server equipped with two NVIDIA A100 80GB PCIe GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03.
+Java SFT hyperparameters:
+For Java experiments, we fine-tune Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct with bfloat16 precision.
+Both models are trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and warmup ratio of 0.1.
+We use a maximum sequence length of 3000 tokens, per-device batch size of 2, and gradient accumulation steps of 4.
+Gradient checkpointing (non-reentrant) is enabled.
+Evaluation and checkpoint saving occur every 5000 steps.
+Special tokens are manually handled in the chat template.
+Training is seeded with 42.
+Java RS-DPO parameters:
+RS-DPO training is performed on Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct using bfloat16 precision.
+Training runs for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and warmup ratio of 0.1.
+We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4.
+Gradient checkpointing (non-reentrant) is enabled.
+Evaluation and checkpoints are recorded every 200 steps.
+Reward shaping parameters vary across settings, with $\beta\in\{0.1,0.5,1\}$.
+Seeds are fixed at 42 for reproducibility.
+Computational Environment:
+All SFT, RS-SFT, and RS-DPO experiments (Python and Java) were conducted on a Linux server equipped with NVIDIA A100 80GB GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03.
 Each job had access to 100 GB of CPU memory and 2 CPU cores.
-Training was performed using mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage.
-For inference we used a similar environment but varied number of GPUs based on model sizes and needs.
-A.1.4 Baseline Inference Details
+Training used mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage.
+Inference used a similar setup with GPU allocation varying by model size.
+C.4 Baseline Inference Details
 We use the following hyperparameters for performing inference with the baseline LLMs:
@@
 We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty).
+For GPT-5 we use 8192 max completion tokens and high reasoning effort.
 Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches.
@@
 Do not detect any idioms not specified above.
-Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called “### Final Idiom Violations Found”.
+Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called ``### Final Idiom Violations Found''.
 Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
 […]
-Inquiry.objects.filter(owner=kwargs.pop(’user’)))", "fix": null}
+Inquiry.objects.filter(owner=kwargs.pop('user')))", "fix": null}
 […]
@@
 and as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:** NO VIOLATIONS FOUND Violations per idiom:
-A.2 Additional Experimental Details
-A.2.1 Idioms Chosen for Ruff Idiom Transfer Dataset
+JEP# JEP Title Definition Example(s) Tree Sitter Queries
+394
+PatternMatching
+InstanceOf
+Usage of the old pattern
+of testing with instanceof
+followed by a manual cast
+to extract and operate on
+the object. This pattern is
+verbose and repetitive. Flag
+the instanceof expression
+check within a conditional
+statement and the
+accompanying cast
+expression in the body of
+the conditional statement.
+public class ShapeExample {
+static double getPerimeter(Object obj) {
+if (obj instanceof Rectangle) {
+Rectangle r = (Rectangle) obj;
+return 2 * r.length() + 2 * r.width();
+} else if (obj instanceof Circle) {
+Circle c = (Circle) obj;
+return 2 * c.radius() * Math.PI;
+} else {
+throw new IllegalArgumentException(
+"Unrecognized shape");
+(if_statement
+condition: (parenthesized_expression
+(instanceof_expression
+left: (identifier) @H1
+right: (type_identifier) @H2
+) @jep_394_before_instanceof_expression.part1
+consequence: (block
+((local_variable_declaration
+type: (type_identifier) @H3
+declarator: (variable_declarator
+value: (cast_expression
+type: (type_identifier) @H4
+value: (identifier) @H5
+)(#eq? @H1 @H5) (#eq? @H2 @H3) (
+#eq? @H3 @H4)
+) @jep_394_before_instanceof_expression.part2
+394
+PatternMatching
+InstanceOf
+Replaces verbose instanceof
+tests plus manual casting into
+a concise form that tests and
+declares a typed variable in
+one step, for example,
+"if (obj instanceof String s)"
+which improves readability,
+reduces boilerplate, and
+introduces flow-scoped pattern
+variables. Flag only the line
+containing the combined
+instancesof test and casting
+within the conditional
+statement.
+public class ShapeExample {
+static double getPerimeter(Object obj) {
+if (obj instanceof Rectangle r) {
+return 2 * r.length() + 2 * r.width();
+} else if (obj instanceof Circle c) {
+return 2 * c.radius() * Math.PI;
+} else {
+throw new IllegalArgumentException(
+"Unrecognized shape");
+(instanceof_expression
+left: (_)
+right: (type_identifier)
+name: (identifier)
+) @jep_394_after_instanceof_expression
+378
+TextBlocks
+Multiline strings represented
+using concatenated string
+literals, requiring explicit
+newline escape sequences
+[…]
+with the + operator. This
+approach is verbose and
+error-prone. Flag cases
+where a variable
+declaration or method
+invocation uses
+concatenated string literals
+instead of multiline strings.
+[…]
+[…]
+[…]
+[…]
+[…]
+(local_variable_declaration
+declarator: (variable_declarator
+name: (identifier)
+value: [
+(binary_expression
+] @jep_378_before_concatenated_string_literals
+378
+TextBlocks
+Use of multiline string literal
+enclosed by triple
+double-quote marks ("""),
+allowing for cleaner and more
+readable representation of
+multiline strings without
+explicit escape sequences.
+Flag cases that use triple
+double-quote marks for
+multiline strings in variable
+declarations or
+method invocations.
+String html = """
+<html>
+<body>
+<p>Hello, world!</p>
+</body>
+</html>
+(string_literal) @jep_378_after_text_block
+(#match?
+@jep_378_after_text_block "^\"\"\"")
+361
+Switch
+Expressions
+Misuse of switch statement
+with fall-through behavior
+'for pattern matching. This
+pattern is verbose and error
+prone. You should flag case
+statements with empty
+bodies that are misusing
+fall-through behavior.
+int numLetters;
+switch (day) {
+case MONDAY:
+case FRIDAY:
+case SUNDAY:
+numLetters = 6;
+break;
+case TUESDAY:
+numLetters = 7;
+break;
+throw new IllegalStateException(
+"Unexpected value: " + day);
+jep_361_before_custom_detectors
+Table 5: JEP Idiom Specifications (1/3): This table presents 15 idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns.
+The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern.
+The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag idioms for synthetic data creation.
+JEP# JEP Title Definition Example(s) Tree Sitter Queries
+361
+Switch
+Expressions
+Use of switch expressions, allowing
+a return value. Employs the ->
+syntax for case labels, eliminating
+fall-through behavior. Flag
+statements that use the arrow
+operator "->" or "yield" syntax.
+Example 1:
+int numLetters = switch (day) {
+case MONDAY, FRIDAY, SUNDAY ->6;
+case TUESDAY ->7;
+case THURSDAY, SATURDAY ->8;
+case WEDNESDAY ->9;
+Example 7:
+String category = switch (age) {
+case 0, 1, 2, 3, 4, 5 ->"Toddler";
+case 6, 7, 8, 9, 10, 11, 12 ->"Child";
+case 13, 14, 15, 16, 17, 18, 19 ->"Teenager";
+default ->"Adult";
+Example 8:
+String response = switch (input) {
+case "yes" ->"Affirmative";
+case "no" ->"Negative";
+default ->"Unrecognized input";
+(yield_statement
+) @jep_361_after_yield
+(switch_rule
+(switch_label)
+"->" @jep_361_after_arrow
+(switch_rule
+(switch_label)
+"->" ;; ensures it's not arrow
+(block (yield_statement
+) @jep_361_after_yield)
+314
+UnicodeLang
+TagExtensions
+Use java.util.Locale with additional
+BCP 47 Unicode extensions
+(cu, fw, rg, tz) in Java 10 to customize
+locale behavior like currency
+(java.util.Currency), first-day-of-week
+(java.time.temporal.WeekFields),
+region override
+(java.text.NumberFormat.getInstance),
+and time zone
+(java.time.format.DateTimeFormatter).
+Flag imports and function calls
+related to these.
+Example 1: Currency Type (cu)
+import java.util.Locale;
+import java.util.Currency;
+public class Foo {
+void bar() {
+Locale locale = Locale.forLanguageTag(
+"en-US-u-cu-EUR");
+Currency c = Currency.getInstance(locale);
+System.out.println(c);
+Example 4: Time Zone (tz)
+import java.util.Locale;
+import java.time.format.DateTimeFormatter;
+import java.time.ZonedDateTime;
+public class Foo {
+void bar() {
+Locale locale = Locale.forLanguageTag(
+"en-US-u-tz-Asia-Tokyo");
+DateTimeFormatter fmt =
+DateTimeFormatter.ofPattern(
+"yyyy-$M$-$d$ $H$:$m$ z").withLocale(locale);
+System.out.println(
+fmt.format(ZonedDateTime.now()));
+(import_declaration
+((scoped_identifier
+scope: (scoped_identifier
+) @H2
+name: (identifier) @H1
+) (#eq? @H2 "java.util") (
+#eq? @H1 "Currency"))
+) @jep_314_after_currency_import
+((method_invocation
+object: (identifier) @H7
+name: (identifier) @H8
+) (#eq? @H7 "Currency") (
+#eq? @H8 "getInstance")
+) @jep_314_after_currency…
+((method_invocation
+object: (identifier) @H13
+name: (identifier) @H14
+#eq? @H13 "NumberFormat") (
+#eq? @H14 "getInstance")
+) @jep_314_after_number_format…
+395
+RecordClass
+Use of simple data aggregates with
+traditional classes which could be
+replaced with a record class. This
+approach requires explicit
+declarations of fields, constructors,
+and accessor methods, leading to
+verbose and repetitive code. Flag
+non record classes containing equals(),
+hashCode(), and toString() methods.
+public class Point {
+private final int x;
+private final int y;
+public int x() {
+return x;
+@Override
+public String toString() {
+return "Point{x=" + x + ", y=" + y + "}";
+@Override
+public boolean equals(Object obj) {
+@Override
+public int hashCode() {
+return Objects.hash(x, y);
+(class_declaration
+body: (class_body
+(constructor_declaration
+) @H1
+(method_declaration
+name: (identifier) @H2
+) (#match? @H2 "^(
+hashCode|equals|toString)\$")
+) @jep_395_before_record_like_class
+395
+RecordClass
+Use of record class. Record classes
+introduce a concise syntax for
+defining immutable data aggregates,
+automatically generating canonical
+constructors, accessors, equals(),
+hashCode(), and toString() methods,
+thereby reducing boilerplate code
+and enhancing readability.
+Example 1 (Record Declaration):
+record Point(int x, int y) {}
+Example 2 (Record Declaration):
+record Rectangle(double length, double width) {}
+(record_declaration
+) @jep_395_after_record_…
+Table 6: JEP Idiom Specifications (2/3)
+JEP# JEP Title Definition Example(s) Tree Sitter Queries
+409
+Sealed
+Class
+Use of abstract classes with
+private constructors to simulate
+sealed classes using
+package-private visibility to
+restrict subclassing. This
+approach lacks explicit language
+support and is error-prone.
+Switch to sealed classes. Flag
+abstract classes with private
+constructors.
+public abstract class Shape {
+private Shape() {}
+public class Circle extends Shape {
+/* Implementation */
+public class Square extends Shape {
+/* Implementation */
+((class_declaration
+(modifiers) @H1
+name: (identifier) @H4
+body: (class_body
+(constructor_declaration
+(modifiers) @H2
+name: (identifier) @H3
+… )(#match? @H1 "abstract")
+(#eq? @H3 @H4)
+(#match? @H2 "private")
+) @jep_409_before_abstract_class…
+409
+Sealed
+Class
+Use of sealed classes to
+explicitly define which classes
+or interfaces can extend or
+implement them using the
+sealed modifier and the permits
+clause. This feature enhances
+type safety and exhaustiveness
+checking. Flag class declarations
+with the sealed or non-sealed
+modifiers and lines with the
+permit clause.
+public sealed class Shape
+permits Circle, Square {
+/* Implementation */
+public final class Circle extends Shape {
+/* Implementation */
+public final class Square extends Shape {
+/* Implementation */
+(permits
+) @jep_409_after_permits_clause
+((class_declaration
+(modifiers) @H1
+)(#match? @H1 "sealed")
+) @jep_409_after_sealed_modifier
+406
+Pattern
+Matching
+Switch
+Use of a sequence of if-else
+if statements to test an object's
+type via instanceof, with a
+manual cast, to handle each case
+separately. This approach is
+verbose, error-prone, and lacks
+exhaustiveness checking or
+compiler assistance for missing
+cases. Flag if or else-if statements
+that contain instanceof statements
+with a manual cast in the
+statement body.
+static String formatter(Object o) {
+if (o instanceof Integer) {
+Integer i = (Integer) o;
+return String.format("int %d", i);
+} else if (o instanceof Long) {
+Long l = (Long) o;
+return String.format("long %d", l);
+} else if (o instanceof String) {
+String s = (String) o;
+return String.format("String %s", s);
+} else {
+return o.toString();
+(if_statement
+condition: (parenthesized_expression
+(instanceof_expression
+left: (identifier) @H1
+right: (type_identifier) @H2
+) @jep_406_before_if_else_if_…
+consequence: (block
+((local_variable_declaration
+type: (type_identifier) @H3
+declarator: (variable_declarator
+value: (cast_expression
+type: (type_identifier) @H4
+value: (identifier) @H5
+(#eq? @H1 @H5)
+(#eq? @H2 @H3)
+(#eq? @H3 @H4)
+) @jep_406_before_if_else_if…
+406
+Pattern
+Matching
+Switch
+Use of a switch expression or
+statement with case labels
+containing type patterns (and
+optionally a guard), binding
+the matched variable within
+the branch. This style is more
+concise, expressive, and opens
+opportunities for compiler-checked
+exhaustiveness and performance
+optimizations. Flag switch labels
+(case statements) with patterns, null
+literals or paranthesized expressions
+but skip default switch labels/cases.
+Example 1:
+static String formatter(Object o) {
+return switch (o) {
+case Integer i ->String.format("int %d", i);
+case Long l ->String.format("long %d", l);
+case String s ->String.format("String %s", s);
+default ->o.toString();
+Example 2:
+static String checkShape(Object o) {
+return switch (o) {
+(switch_label
+(null_literal)
+) @jep_406_after_null_case
+(switch_label
+) @jep_406_after_switch_pattern
+(switch_label
+(parenthesized_expression)
+) @jep_406_after_paranthesized_pattern
+(switch_label
+(binary_expression)
+) @jep_406_after_binary_expression
+323
+LocalVar
+Syntax
+Lambda
+Params
+Use of implicitly typed lambda
+expressions with omitted type
+declarations. These lambda
+expressions rely solely on parameter
+names. This approach prioritizes
+brevity but lacks explicit type
+information. Flag full lambda
+expressions without type
+declarations.
+Example 1:
+xs.stream().filter((a, b) ->a <b).forEach(
+System.out::println);
+Example 4:
+xs.stream().filter((a) ->a >10).forEach(
+System.out::println);
+jep_323_before_custom_detector
+323
+LocalVar
+Syntax
+Lambda
+Params
+Use of explicit type declarations for
+lambda parameters, enhancing code
+clarity and enabling better static
+analysis tools. Flag full lambda
+expressions with explicit type
+declarations using formal
+parameters (var).
+Example 1:
+xs.stream().filter(
+(var a, var b) ->a.compareTo(b) <0).forEach(
+System.out::println);
+Example 4:
+xs.stream().filter((var a) ->a >10).forEach(
+System.out::println);
+(lambda_expression
+parameters: (formal_parameters
+(formal_parameter
+type: (type_identifier) @H1
+(#eq? @H1 "var"))
+)) @jep_323_after_local_var_lambda
+Table 7: JEP Idiom Specifications (3/3)
+C.5 PMD Idiom Specifications
+We scrape PMD idioms specification from the Java section of the PMD rules documentationhttps://docs.pmd-code.org/latest/pmd_rules_java.html.
+The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as idiom violations, as can be seen in the example below.
+PMD Rule Specification: UnitTestShouldIncludeAssert
+Since: PMD 2.0
+Priority: Medium (3)
+Unit tests should include at least one assertion. This makes
+the tests more robust, and using assert with messages provide
+the developer a clearer idea of what the test does. This rule
+checks for JUnit (3, 4 and 5) and TestNG Tests. Note: This rule
+was named JUnitTestsShouldIncludeAssert before PMD 7.7.0. This
+rule is defined by the following Java class:
+net.sourceforge.pmd.lang.java.rule.bestpractices.
+UnitTestShouldIncludeAssertRule
+Example(s):
+public class Foo {
+   @Test
+   public void testSomething() {
+      Bar b = findBar();
+      // This is better than having a NullPointerException
+      // assertNotNull("bar not found", b);
+      b.work();
+This rule has the following properties:
+Name
+Default Value
+Description
+extraAssertMethodNames
+Extra valid assertion methods names
+Use this rule with the default properties by just referencing
+it:
+<rule ref="category/java/bestpractices.xml/
+UnitTestShouldIncludeAssert" />
+Use this rule and customize it:
+<rule ref="category/java/bestpractices.xml/
+UnitTestShouldIncludeAssert">
+    <properties>
+        <property name="extraAssertMethodNames" value="" />
+    </properties>
+</rule>
+JEP # Before After Title JDK# Release Date
+409 Yes Yes Sealed Classes 17
+406 Yes Yes Pattern Matching for switch 17 14 Sept 2021
+395 Yes Yes Records 16
+394 Yes Yes Pattern Matching for instanceof 16 16 Mar 2021
+378 Yes Yes Text Blocks 15 15 Sept 2020
+361 Yes Yes Switch Expressions 14 17 Mar 2020
+323 Yes Yes
+Local-Variable Syntax for
+Lambda Parameters
+11 25 Sept 2018
+314 No Yes
+Additional Unicode
+Language-Tag Extensions
+10 20 Mar 2018
+Table 8: List of JEPs addressed by our tree-sitter synthetic data.
+The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features.
+The Before and After columns indicate whether we include rules/patterns to flag the old idiom or new idiom introduced by the JEP.
+Java MetaLint Instruction Following Prompt Task Instructions (1/2): Look at the following code idiom specification with definitions and examples: {IDIOM_SPEC} Task Instructions (2/2): Given this idiom, your task is to look at a code file and detect violations of the above idiom, and flag them like a linter.
+You should also suggest a fix if possible.
+Report the results for only the idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for the given idiom.
+Do not detect violations of any idiom not specified above.
+Code file: {CODE_FILE} Violations per idiom:
+Appendix D Additional Experimental Details
+D.1 Evaluation Metrics
+Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations.
+For each dataset instance with input prompt $x$ and ground truth set of line numbers $y$, $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$.
+We define the indicator variable:
+$\mathds{1}[x]=\begin{cases}1&\text{if }x\text{ is true}\\ 0&\text{otherwise}\end{cases}$
+Detection Metrics:
+$P_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|=0]\cdot\mathds{1}[|\hat{y}|>0]\right)}$
+$R_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|=0]\right)}$
+Macro-averaged detection metrics:
+$P_{\text{Det}}=\frac{1}{|I|}\sum_{I}P_{I},\quad R_{\text{Det}}=\frac{1}{|I|}\sum_{I}R_{I},\quad F_{\text{Det}}=\frac{2P_{\text{Det}}R_{\text{Det}}}{P_{\text{Det}}+R_{\text{Det}}}$
+Localization Metrics:
+$P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
+D.2 Idioms Chosen for Ruff Idiom Transfer Dataset
 Table 9 lists the Ruff idioms used in the SFT training and synthetic transfer evaluation test sets.
@@
 Figure 3: ID: In-Domain, NeT: Near Transfer, FaT: Far Transfer.
-A.2.2 DPO No Violation Fraction Ablations
+D.3 DPO No Violation Fraction Ablations
 We analyze the impact of varying the amount of samples with zero violations used for RS-DPO training.
@@
 To investigate this trade-off, we experimented with keeping only a fraction of the NO VIOLATIONS FOUND data during training.
-Specifically, we randomly sampled $k$of such data, varying $k$ across {0%, 2%, 5%, 10%, 20%, 40%, 100%}.
+Specifically, we randomly sampled kk% of such data, varying $k$ across {0%, 2%, 5%, 10%, 20%, 40%, 100%}.
 These percentages were selected based on observed trends: 20%, 40%, and 100% yielded similar results, which discouraged further tests at 60% or 80%, while 2% and 5% were chosen due to a noticeable performance jump between 0% and 10%.
@@
 Based on these insights, we conducted a limited ablation on the CoT model, evaluating 2% and 5% inclusion to determine the optimal setting for both detection and localization (as shown in Table 11).
-Fraction of NV data Det. Line overlap (Det. + Loc.)
-P R F P R F
+Fraction of
+NV data
+Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
 0% 0.6268 0.9577 0.7577 0.6777 0.6932 0.6854
@@
 All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Idiom Transfer test set.
-Fraction of NV data Det. Line overlap (Det. + Loc.)
-P R F P R F
+Fraction of
+NV data
+Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
 2% 0.9226 0.8901 0.906 0.7688 0.7638 0.7663
@@
 We perform limited ablations because of the insights from the non CoT model training.
-A.3 More Results
-A.3.1 Expanded Results on the PEP Hard Idiom Benchmark
+Fraction of
+NV data
+Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+1% 0.654 0.6468 0.6504 0.491 0.4788 0.4848
+2% 0.6636 0.6057 0.6333 0.4869 0.4745 0.4806
+Table 12: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Llama3.2-3B-Instruct model.
+We perform limited ablations because of the insights from the non CoT model training.
+D.4 PEP Benchmark Creation Additional Details
+As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP idioms. These are summarized in Table 13, 14 and 15.
+PEP Description Heuristics Example
+506
+Adds secrets module
+to the standard library
+for cryptographically secure
+random value generation
+Conjunction of 2 conditions:
+1. Presence of "random" module imports
+2. Presence of "random" function usage
+characters = string.ascii_letters +
+string.punctuation + string.digits
+password = "".join(random.choice
+(characters) for x in range(16))
+Use instead:
+characters = string.ascii_letters +
+string.punctuation + string.digits
+password = "".join(secrets.choice
+(characters) for x in range(16))
+557
+Introduces the dataclasses
+module, enabling automatic
+generation of common boilerplate
+methods for classes
+Conjunction of 2 conditions:
+1. There is a class with manual
+implementation of "__init__" method
+2. On the same class there is manual
+implementation of common
+special methods or comparison methods
+that follow standard data storage patterns.
+"class Point:
+def __init__(self, x, y):
+self.x = x
+self.y = y
+def __repr__(self):
+return f""Point(x={self.x}, y={self.y})""
+Use instead:
+from dataclasses import dataclass
+@dataclass
+class Point:
+x: int
+y: int"
+655
+Introduces Required[] and
+NotRequired[] type qualifiers
+to replaces cumbersome
+TypedDict inheritance patterns.
+Conjuction of:
+1. "TypedDict" defined with inheritance pattern.
+2. total=False parameter usage in class definition
+class _MovieBase(TypedDict): # implicitly total=True
+title: str
+class Movie(_MovieBase, total=False):
+year: int
+Use instead:
+class Movie(TypedDict):
+title: str
+year: NotRequired[int]
+634
+Introduced structural pattern
+matching, enabling more
+expressive and concise ways
+to match data structures
+and control flow.
+Multiple consecutive if-elif-else
+statements that compare a
+single variable against different
+values with dysjunction of 2 conditions:
+1. Length of ladder (number of
+conditons at the "top level" + one level in) >= 6
+2. Depth of ladder (degree of nesting) >=3
+"def handle_response(response):
+if isinstance(response, dict):
+if ""error"" in response:
+print(f""Error: {response['error']}"")
+elif ""data"" in response:
+print(f""Data: {response['data']}"")
+else:
+print(""Unknown response format"")
+elif isinstance(response, list):
+print(""List of items:"", response)
+else:
+print(""Invalid response type"")
+Use instead:
+def handle_response(response):
+match response:
+case {""error"": error_message}:
+print(f""Error: {error_message}"")
+case {""data"": data_content}:
+print(f""Data: {data_content}"")
+case list(items):
+print(""List of items:"", items)
+case _: print(""Invalid response type"")"
+614
+Removes previous restrictions
+on decorator syntax. Before,
+only simple names or dotted
+names were valid decorators.
+After 614, any valid expression
+can be used as a decorator
+Conjunction of 2 conditions:
+1. A decorator is applied using a name
+(e.g., @decorator) where that name is
+assigned earlier in the code.
+2. The assignment value is an expression
+of type Call, Attribute, or Subscript (e.g.,
+deco = factory(), deco = module.decorator,
+deco = decorators[i]).
+# def uppercase(func):
+def wrapper(*args, **kwargs):
+return func(*args, **kwargs).upper()
+return wrapper
+@uppercase
+def greet():
+return "hello"
+Use Instead:
+deco = [uppercase]
+@deco[0]
+def greet2():
+return "hi"
+616
+Replaces manual slicing
+with dedicated methods
+dysjunction of 2 conditions:
+1. There is a "check" with
+startswith or endswith on a
+given variable x.
+2. On the same variable x
+check if there is an "edit" using a
+program slicing syntax or using "replace()".
+if s.startswith(prefix): s = s[len(prefix):]
+Use instead:
+s = s.removeprefix(prefix)
+OR
+s[:-len(suffix)]
+Use instead:
+s.removesuffix(suffix)
+584
+Introduces the binary operators
+| (merge) and |= (update) on
+dict (and other built-in mapping
+types), providing an expressive,
+in-place-or-new‐object way to
+combine dictionaries.
+disjunction of two conditions:
+1. A copy-and-update sequence
+on the same variable
+or in close proximity: d = d1.copy()
+followed by d.update(d2)
+2. A dictionary literal using multiple
+unpackings {**d1, **d2},
+indicating ad-hoc merging rather
+than the new operators
+d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
+merged = d1.copy()
+merged.update(d2)
+d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
+merged = {**d1, **d2}
+d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
+merged = dict(list(d1.items()) + list(d2.items()))
+Use instead:
+d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
+merged = d1 | d2
+d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
+d1 |= d2 # d1 is now {'a': 1, 'b': 2, 'c': 3, 'd': 4}
+Table 13: High recall heuristics used to find instances of PEP violations that human annotators vet
+PEP Description Heuristics Example
+570
+Introduces new syntax
+(the / marker) in Python
+function signatures to
+specify positional-only
+parameters, ensuring that
+certain arguments can
+only be supplied by their
+position and not as keywords
+Conjunction of the
+following conditions:
+1. Have only positional-or-
+keyword parameters
+(without *args, **kwargs, keyword
+-only parameters, or the '/' marker),
+2. Include 2 to 4 parameters, all of
+which have no default values
+def compute_area(width, height):
+return width * height
+area = compute_area(width=5, height=10)
+print("Area:", area)
+Use instead:
+def compute_area(width, height, /):
+return width * height
+area = compute_area(5, 10)
+print("Area:", area)
+567
+Adds the contextvars module,
+enabling context-local variables
+for managing dynamic state.
+Dysjunction of the following conditions:
+1. Look for import threading together
+with threading.local() object creation and use.
+2. Find global statements or assignment
+to variables at the module
+level that are accessed or mutated in functions,
+especially as shared state.
+3. Identify async functions or classes
+where context or state
+variables are passed as parameters
+(e.g., def func(context, …) or async
+def func(context, …)), not as context-local variables.
+import threading
+_thread_local = threading.local()
+def set_context(value):
+_thread_local.value = value
+def get_context():
+return getattr(_thread_local, 'value', None)
+Use instead:
+from contextvars import ContextVar
+context_var = ContextVar('value')
+def set_context(value):
+context_var.set(value)
+def get_context():
+return context_var.get()
+530
+Enables the use of "async for"
+and "await" in list, set, and dict
+comprehensions as well as in
+generator expressions, providing
+concise asynchronous data
+processing within comprehensions
+Dysjunction of the following conditions:
+1. "async" def functions that uses "async for"
+loops to build lists, sets, or dicts.
+2. "async for" loops, followed by methods like
+result.append(…), result.extend(…), or result[key] = ….
+3. Comprehensions written without the "async for"
+clause despite being inside an "async def"
+result = []
+async for i in aiter():
+if i % 2:
+result.append(i)
+Use instead:
+result = [i async for i in aiter() if i % 2]
+525
+Introduces the ability to define
+asynchronous generator functions
+using the async def and yield
+syntax, enabling concise, native
+support for asynchronous iteration.
+Dysjunction of the following conditions:
+1. classes defining both "__aiter__" and "__anext__"
+methods, especially where the class is used solely to
+produce a sequence of values asynchronously.
+2. async def functions that create and return custom
+iterator classes instead of using async def with yield.
+class Ticker:
+"""Yield numbers from 0 to `to`
+every `delay` seconds."""
+def __init__(self, delay, to):
+self.delay = delay
+self.i = 0
+self.to = to
+def __aiter__(self):
+return self
+async def __anext__(self):
+i = self.i
+if i >= self.to:
+raise StopAsyncIteration
+self.i += 1
+if i:
+await asyncio.sleep(self.delay)
+return i
+Use instead:
+async def ticker(delay, to):
+"""Yield numbers from 0 to `to`
+every `delay` seconds."""
+for i in range(to):
+yield i
+await asyncio.sleep(delay)
+520
+Ensures that the order in
+which attributes are defined within
+a class body is preserved in the
+resulting class object, making
+the attribute order predictable
+and consistent.
+Dysjunction of the following conditions:
+1. Uses sorted() or otherwise processes
+class.__dict__.keys() to impose attribute order.
+2. Attribute names are tracked in a list or similar
+structure solely to maintain definition order.
+3. Custom metaclass logic or "__prepare__" impleme-
+ntations created to preserve the order of class attributes.
+class Person:
+name = "Alice"
+age = 30
+city = "Wonderland"
+def display_attributes(self):
+# Manually sorting keys
+for key in sorted(self.__class__.__dict__.keys()):
+if not key.startswith("__"):
+print(key, getattr(self, key))
+Use instead:
+class Person:
+name = "Alice"
+age = 30
+city = "Wonderland"
+def display_attributes(self):
+# Directly iterate over the preserved definition order
+for key in self.__class__.__definition_order__:
+print(key, getattr(self, key))
+498
+Introduces f-strings (formatted
+string literals) as a new,
+concise, and efficient way
+to embed Python
+expressions inside string
+literals using the f'' prefix.
+Dysjunction of the following conditions:
+1. Occurrences of string literals with
+.format(…) applied,
+especially where keys or variables match
+braces in the string
+2. String literals concatenated using
+"+" with variables.
+3. Uses of the "%" operator for string formatting,
+name = "Alice"
+age = 30
+greeting = "Hello, " + name +
+"! You are " + str(age) + " years old."
+Use instead:
+name = "Alice"
+age = 30
+greeting = f"Hello, {name}!
+You are {age} years old."
+OR
+value = 12.3456
+formatted = "The value is {:.2f}".format(value)
+Use instead:
+value = 12.3456
+formatted = f"The value is {value:.2f}"
+Table 14: High recall heuristics used to find instances of PEP violations that human annotators vet
+PEP Description Heuristics Example
+487
+Makes customizing class
+creation and subclass initialization
+easier by introducing __init_
+subclass__ and __set_name__,
+eliminating the need for
+most custom metaclasses
+Disjunction of the following conditions:
+1. Custom metaclasses defined to
+execute code during class creation
+or subclassing (e.g., overriding __new__,
+__init__, or __call__ in metaclasses)
+instead of using __init_subclass__.
+2. Descriptor classes lacking __set_
+name__ method and employing
+manual workarounds to determine
+their assigned attribute names.
+3. Classes or frameworks manually
+tracking or registering sub-
+classes via metaclass hooks instead
+of leveraging __init_subclass__.
+class Meta(type):
+def __new__(meta, name, bases, namespace):
+for key, value in namespace.items():
+if isinstance(value, Descriptor):
+value.name = key
+return super().__new__
+(meta, name, bases, namespace)
+class MyClass(metaclass=Meta):
+attr = Descriptor()
+Use instead:
+class Descriptor:
+def __set_name__(self, owner, name):
+self.name = name
+class MyClass:
+attr = Descriptor()
+OR
+class PluginBase(type):
+plugins = {}
+def __new__(meta, name, bases, namespace):
+if name != 'Plugin':
+meta.plugins[name] = namespace['priority']
+return super().__new__
+(meta, name, bases, namespace)
+class Plugin(metaclass=PluginBase):
+priority = 0
+class HighPriority(Plugin):
+priority = 10
+Use instead:
+class Plugin:
+plugins = {}
+priority = 0
+def __init_subclass__(cls, **kwargs):
+super().__init_subclass__(**kwargs)
+cls.plugins[cls.__name__] = cls.priority
+class HighPriority(Plugin):
+priority = 10
+593
+Introduces flexible function
+and variable annotations via
+typing.Annotated, which lets you
+attach context-specific metadata
+to type hints (e.g., validation
+constraints, units)
+Conjunction of 2 conditions:
+1. Type hints are already present
+in function arguments, return
+types, or variable annotations.
+2. Nearby comments/docstrings (within
+±2 lines) contain metadata-like
+patterns such as "min", "max",
+"nullable", "regex", "enum",
+"unit", "deprecated", etc.
+# max 100, min 1
+def set_age(age: int) ->None:
+pass
+Use instead:
+from typing import Annotated
+Age = Annotated[int, "min=1", "max=100"]
+def set_age(age: Age) ->None:
+pass
+526
+introduces explicit variable
+annotations, allowing type
+hints directly on variable
+declarations for local, global,
+and class variables in Python
+Disjunction of the following conditions:
+1. Variables assigned values with a
+type comment (e.g., x = 0 # type: int)
+instead of using annotation syntax.
+2. Identify variable assignments,
+especially class and
+instance attributes, that lack any
+type annotation (e.g., name = "" in class bodies).
+3.Module-level variables assigned
+values without accompanying type hints—
+especially in type-annotated codebases.
+# type: List[int]
+numbers = []
+Use instead:
+numbers: List[int] = []
+OR
+class Player:
+# type: str<br>name = "Guest"
+Use instead:
+class Player:
+name: str = "Guest"
+589
+Introduces TypedDict, enabling
+precise type hints for
+dictionaries with a fixed
+set of string keys, improving
+static type checking and
+readability in Python code.
+Disjunction of the following conditions:
+1. Dictionary literals or variables
+consistently using the same fixed set of
+string keys without accompanying
+TypedDict annotations.
+2. Functions annotated with broad
+dictionary types like Dict[str, Any], dict,
+or untyped parameters/returns that actually
+expect dictionaries with a known fixed set of keys.
+3. Explicit key presence checks or
+accessing dictionary keys repeatedly that suggest
+a structured dictionary shape.
+movie = {'name': 'Blade Runner',
+'year': 1982}
+Use instead:
+from typing import TypedDict
+class Movie(TypedDict):
+name: str
+year: int
+movie: Movie = {'name': 'Blade Runner',
+'year': 1982}
+572
+Introduces the assignment
+expression operator :=
+(the "walrus operator"),
+allowing assignment to
+variables within expressions,
+1. Patterns where a value is first assigned
+to a variable, and then immediately checked
+or used in the next line or inside a
+loop, list comprehension, or condition.
+2. separate assignment and conditional test statements
+match = pattern.search(data)
+if match is not None:
+process(match)
+Use instead:
+if (match := pattern.search(data)) is not None:
+process(match)
+Table 15: High recall heuristics used to find instances of PEP violations that human annotators vet
+Appendix E More Results
+Model In-Domain Near Transfer Far Transfer
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
+Qwen3-4B 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
++SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
++RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
+Qwen3-4B w CoT 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
++RS-SFT 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
++RS-DPO 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
+Llama3.2-3B-Instruct 0.54 0.43 0.48 0.69 0.68 0.69 0.47 0.51 0.49
++SFT 0.88 (+0.34) 0.87 (+0.44) 0.88 (+0.4) 0.89 (+0.2) 0.44 (-0.24) 0.59 (-0.1) 0.61 (+0.14) 0.27 (-0.24) 0.37 (-0.12)
++RS-DPO 0.75 (+0.21) 0.92 (+0.49) 0.83 (+0.35) 0.81 (+0.12) 0.71 (+0.03) 0.76 (+0.07) 0.61 (+0.14) 0.59 (+0.08) 0.60 (+0.11)
+Table 16: Cross-Idiom Generalization on Python Ruff Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff idioms, and the performance is reported on other Ruff idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
+E.1 Expanded Results on the PEP Hard Idiom Benchmark
 We show the expanded results across various model sizes for the evaluated model families in Table 17.
@@
 Detection Localization
-Model P R F P R F
+Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Llama3.2-3B-Instruct 0.7042 0.214 0.3283 0.0691 0.0798 0.0741
 Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
@@
 DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
-MetaLint (SFT) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
-MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
-MetaLint w CoT (RS-SFT) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
-Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+GPT-oss-20b 0.8377 0.3531 0.4968 0.251 0.1695 0.2024
+GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
+Qwen3-4B MetaLint (SFT) (Ours) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
+Qwen3-4B MetaLint (SFT+RS-DPO) (Ours) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
+Qwen3-4B MetaLint w CoT (RS-SFT) (Ours) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
+Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) (Ours) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+Llama3.2-3B-Instruct MetaLint (SFT) (Ours) 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
+Llama3.2-3B-Instruct MetaLint (SFT+RS-DPO) (Ours) 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
+o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
+o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
 GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
 GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
-o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
-o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
-Table 7: Result on the hard PEP benchmark to measure easy to hard generalization.
-A.3.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
+GPT-5 (high) 0.913 0.5673 0.6998 0.4397 0.4257 0.4326
+Table 17: Results on the hard PEP benchmark to measure easy to hard generalization.
+Model Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
+Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
+Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
+Qwen3-4B w CoT 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
+Qwen3-4B w CoT + RS-SFT 0.935 0.8183 0.8727 0.6639 0.65 0.6569
+Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.771 0.7571 0.764
+Llama3.2-3B-Instruct 0.5092 0.5286 0.5187 0.1371 0.3 0.1882
+Llama3.2-3B-Instruct + SFT 0.6793 0.3598 0.4704 0.3424 0.3485 0.3454
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.654 0.6468 0.6504 0.491 0.4788 0.4848
+Table 18: Cross-Idiom Generalization on Python Ruff Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff idioms and tested on other Ruff idioms with varying levels of transfer (section 4.2).
+Best score across the compared training setups per model are bolded.
+Model Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
+Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
+Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
+Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
+Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
+Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+Llama3.2-3B-Instruct 0.7042 0.214 0.3283 0.0691 0.0798 0.0741
+Llama3.2-3B-Instruct + SFT 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
+Table 19: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3).
+Best score across the compared training setups per model are bolded.
+E.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
 To analyze the statistical significance of performance differences over the PEP benchmark, we conduct Wilcoxon signed-rank tests comparing various MetaLint variants against each other and against baseline models.
@@
 Table 26: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
-A.3.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
+E.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
 We observe that a significant portion of the lower detection recall of the CoT MetaLint Qwen3-4B model, relative to its non CoT counterpart, can be attributed to its higher tendency to predict NO VIOLATIONS FOUND in cases that do, in fact, contain violations.
@@
 We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__.
-The CoT model often misclassifies such “simple” use cases as complex.
+The CoT model often misclassifies such ``simple'' use cases as complex.
 Overall, these patterns suggest that the CoT model applies the idiom specifications more conservatively, resulting in higher precision but at the cost of reduced recall.
@@
 … the register is a partial function, which is a simple expression. …
-… the decorator is just a variable that’s a partial function. …
+… the decorator is just a variable that's a partial function. …
 Ambiguity about what constitutes
@@
 … The user … wants me to check for … idiom, 525, which
-is about using asynchronous generators with ”async for”
+is about using asynchronous generators with "async for"
 syntax … the violation would be if the code uses
@@
 asynchronous ones. … Looking through the code, I
-don’t see any for loops that are using ’async for’. …
+don't see any for loops that are using 'async for'. …
 CoT accidentally looks for the
@@
 for simple customization tasks makes code unnecessarily
-complex … there’s a class called FirestoreClientMeta,
+complex … there's a class called FirestoreClientMeta,
 which is a metaclass … used for registration of transports
@@
 tasks that could have been handled
-with ”__init_subclass__” or
-”__set_name__”)
+with "__init_subclass__" or
+"__set_name__")
 Table 27: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.
+Model Transfer Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
+Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
+Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.28
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.593 0.5969 0.5949
+Llama3.1-8B-Instruct PMD $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
+Llama3.1-8B-Instruct + SFT 0.2264 0.4508 0.3014 0.3201 0.3152 0.3177
+Llama3.1-8B-Instruct + SFT + RS-DPO 0.4427 0.9191 0.5976 0.6506 0.6709 0.6606
+Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
+Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
+Llama3.1-8B-Instruct PMD $\rightarrow$ JEP 0 0 0 0 0 0
+Llama3.1-8B-Instruct + SFT 0.2166 0.3724 0.2739 0.1455 0.1142 0.128
+Llama3.1-8B-Instruct + SFT + RS-DPO 0.4964 0.8047 0.614 0.4615 0.3395 0.3912
+Llama3.2-3B-Instruct JEP $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
+Llama3.2-3B-Instruct + SFT 0.9567 0.8411 0.8952 0.7837 0.754 0.7686
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.9406 0.86 0.8985 0.7859 0.7651 0.7753
+Llama3.1-8B-Instruct JEP $\rightarrow$ JEP 0 0 0 0 0 0
+Llama3.1-8B-Instruct + SFT 0.9658 0.8466 0.9023 0.809 0.7844 0.7965
+Llama3.1-8B-Instruct + SFT + RS-DPO 0.9308 0.8686 0.8986 0.8131 0.7756 0.7939
+Llama3.2-3B-Instruct JEP $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
+Llama3.2-3B-Instruct + SFT 0.3722 0.2708 0.3152 0.0574 0.0869 0.0692
+Llama3.2-3B-Instruct + SFT + RS-DPO 0.4322 0.4054 0.4183 0.0878 0.1222 0.1022
+Llama3.1-8B-Instruct JEP $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
+Llama3.1-8B-Instruct + SFT 0.3514 0.2229 0.2728 0.0383 0.0753 0.0508
+Llama3.1-8B-Instruct + SFT + RS-DPO 0.436 0.4898 0.4613 0.0831 0.1351 0.1029
+Table 28: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28).
+The transfer column indicates training and test data on the left and right side of the arrow.
+Best score across the compared training setups per model are bolded.

### v3

@@
 Abstract
-Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices.
-We introduce MetaLint, an instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications.
-Unlike conventional approaches that train models on static code quality conventions, MetaLint employs instruction tuning on synthetic linter-generated data with dynamic conventions to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining.
-To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize.
-Our results show that MetaLint training improves generalization to unseen idioms.
-Qwen3-4B attains a 70.37% F-score on a manually curated and challenging PEP idiom detection benchmark, achieving the highest recall (70.43%) among all evaluated models.
-For localization, it reaches 26.73%, which is a strong outcome for its 4B parameter size and comparable to larger state-of-the-art models such as o3-mini, highlighting its potential for future-proof code quality analysis.
-Furthermore, MetaLint training enables generalization in idiom detection across model families, model scales, synthetic data from diverse linters, and Java idioms, demonstrating the general applicability of our approach.
-We plan to release our code and data to enable reproducibility and further work.
+Large Language Models excel at code generation but struggle with code quality analysis, where best practices evolve and cannot be fully captured by static training data.
+We introduce MetaLint, a training framework that treats code quality analysis as detecting best practice violations from high-level specifications over semantic code fragments (code idioms).
+Instead of training on a fixed set of rules, MetaLint reorganizes supervision around dynamically specified best practices using synthetic linter-derived labels, integrated with instruction-following and preference optimization.
+This encourages extrapolation to more complex, unseen best practices at test time, consistent with easy-to-hard generalization without retraining.
+To evaluate MetaLint, we create a new benchmark of hard-to-detect best practices inspired by Python Enhancement Proposals.
+Across this benchmark, MetaLint improves generalization to unseen best practices.
+Qwen3-4B achieves a 2.7× detection F-score gain (25.9% → 70.4%), the highest recall, and a 26.7% localization F-score, matching larger models such as o3-mini.
+These gains generalize across programming languages, model families, scales, reasoning settings, and linter sources.
+Easy to Hard Generalization, Transfer Learning
 1 Introduction
-With the rise of Large Language Models (LLM) of code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent Singhal et al.
-(2024); Zheng et al.
-(2024).
-Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Singhal et al., 2024; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023).
-Several evaluation studies indicate that LLMs struggle with this task Singhal et al.
-(2024); Zheng et al.
-(2024), while attempts to improve them through prompting or training are limited by task-specific, static datasets often grounded in narrow or outdated coding practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b).
-As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024).
-They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024).
-Ideally, we would develop LLM systems that can identify code quality issues without explicit supervision for target idioms—especially hard or rare patterns—and adapt to evolving best practices over time.
-We approach this problem by training the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms.
-For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function.
-However, as noted in PEP 506 (D'Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against using this module for security reasons, which is often missed by developers, as highlighted by accepted answers on forums like StackOverflow.
-PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom.
-As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024).
-Additionally, these issues can be addressed by replacing instances of ``bad'' idioms with corresponding ``good'' idioms that align with best practices.
-Moreover, for this example and similar abstract idioms, constructing a precise rule-based approach is difficult.
-Simply flagging any use of random.choice, even in non-security-critical scenarios (e.g., randomization in a game engine), could result in a poor user experience.
-Vijayvergiya et al.
-(2024) show that LLMs can capture abstract notions of code quality, such as code idioms where building a linter or rule-based approach is challenging, by incorporating semantic reasoning about code and developer intent.
-In this work, we train LLMs to recognize code idioms through a higher-level instruction-following task dubbed ``meta-linting'': given a specification of a best-practice code idiom $I$, the model learns to identify and localize non-idiomatic code fragments.
-Our pipeline is designed to support easy-to-hard generalization (Sun et al., 2024b).
-The easy cases involve simple idioms that can already be captured by existing linters, while the hard cases correspond to nuanced patterns such as PEP 506, where constructing precise rule-based checks is infeasible.
-To enable this, we generate synthetic training data for easy idioms using available linters and leverage it to improve performance on harder cases where linter support is lacking.
-While prior work such as Zhang et al.
-(2024c; b) has explored automated refactoring of non-idiomatic Python code, including the use of LLMs with prompting, our focus differs in three ways.
-First, we target challenging idioms beyond the reach of current linters.
-Second, we train on easy idioms with the goal of transferring detection ability to harder cases.
-Finally, we emphasize adaptability, aiming for LLMs that can accommodate evolving best practices provided in-context as instructions and examples, rather than memorizing a static rule sets.
-To tackle meta-linting, we introduce MetaLint, a training framework motivated by prior work showing that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022).
-Since meta-linting treats each idiom as a distinct task or code quality judgment, instruction fine-tuning (IFT) and preference optimization (PO) naturally extend detection ability to novel idioms.
-Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing simple idioms, which we use both for supervised IFT and as verifiers during PO to improve performance on harder idioms.
-To systematically study this generalization, we construct a benchmark of challenging idioms derived from popular PEPs introducing high-level constructs.
-We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy idioms.
+Code quality analysis offers a unique opportunity to study generalization in Large Language Models (LLMs) because best practices are contextual, evolving, and poorly captured by static training data.
+With the rise of LLMs for code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent (Singhal et al., 2024; Zheng et al., 2024).
+Researchers have explored LLMs’ potential to improve code quality through benchmarks, code review agents, and static analysis (Singhal et al., 2024; Vijayvergiya et al., 2024; Fang et al., 2025).
+However, studies show that LLMs often struggle with these tasks, and improvements via prompting or training are limited by static datasets reflecting narrow or outdated best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b).
+As a result, these systems often perform poorly on rare best practices, different code distributions (Holden & Kahani, 2024) or over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024).
+This difficulty stems not only from limited or outdated training data, but also from supervision: models are trained on a fixed set of best-practice labels, which encourages memorization of individual rules rather than learning reusable abstractions.
+More broadly, generalization to harder or novel violations depends less on model scale or optimization and more on how the learning problem is structured (Chang & Bisk, 2025).
+Motivated by this view, we frame code quality analysis as meta-learning (or ``meta-linting'') over code idioms, recurring patterns that capture semantic intent beyond surface syntax.
+In this formulation, best practices are semantic patterns over idioms, and violations occur when problematic idioms appear in contexts where safer or more appropriate alternatives exist (see Appendix C).
+We operationalize this by using linters to generate synthetic supervision for idioms, then train LLMs to detect and locate violations from high-level natural language specifications and illustrative examples (see Figure 1).
+Code idioms provide a natural intermediate abstraction, bridging simple syntactic cues and full program semantics.
+This allows models to leverage simpler idioms to support detection of more complex, context-dependent best practices.
+For example, in Python, generating passwords or tokens using random.choice is common but insecure in security-sensitive contexts.
+PEP 506 (D'Aprano, 2017) recommends using secrets.choice instead.
+Detecting such violations requires conditional reasoning: not every random.choice is problematic, but its use in security-critical code constitutes a violation.
+Our approach trains models to recognize both the idiom (random.choice) and the relevant context, allowing them to flag violations that go beyond simple syntactic patterns.
+This highlights a broader learning challenge.
+Some best practices are relatively easy to detect because they rely on surface-level cues.
+For example, existing linters such as Ruff include rules (e.g., S104–S108) that flag likely hard-coded secrets using heuristic signals like credential-shaped strings or suspicious variable names.
+While limited and incomplete, these rules expose recurring cues associated with security-sensitive code.
+Meanwhile, hard best practices, such as those captured by PEP 506, build on these same cues but require additional semantic and contextual reasoning.
+Reliably detecting insecure randomness involves recognizing when such signals co-occur with developer intent, such as credential generation or authentication logic, going beyond what fixed heuristics can capture.
+This defines a difficulty progression in code quality analysis driven by the increasing contextual and semantic reasoning required to recognize best practice violations.
+As observed by Chang & Bisk (2025), failures to generalize along such progressions stem not from data scarcity, model scale, or inference limits, but from learning paradigms that fix the hypothesis class and inhibit abstraction.
+In code quality analysis, flat supervision maps programs directly to violation labels, forcing models to memorize isolated best practices rather than extract reusable structure.
+Although easy violations expose low-level heuristics, current training setups do not allow these signals to be elevated into higher-order representations, preventing inductive transfer to harder or novel cases.
+In contrast, MetaLint restructures supervision around code idioms, using signals from easy violations to learn semantic abstractions that become the basis for detecting harder best practice violations from high-level specifications.
+MetaLint integrates naturally with existing instruction fine-tuning (IFT) and preference-optimization (PO) methods, using them for exploiting task structure to achieve cross-task generalization (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022).
+Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing easy best practices, which we use both for supervised IFT and as verifiers during PO to improve performance on hard best practices.
+To systematically study this generalization, we manually construct a benchmark of hard best practices derived from popular PEPs requiring contextual reasoning.
+We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy heuristics.
 Our key contributions are:
   1. 1.
-    We introduce MetaLint, a training framework that leverages instruction following and synthetic data to enable easy-to-hard generalization while remaining adaptable to evolving best practices.
+    We introduce MetaLint, a training framework that reorganizes supervision to bootstrap learning signals for easy best practices from linters to support inductive transfer to hard and novel best practices.
   2. 2.
-    We construct a benchmark of challenging, broadly relevant code-quality idioms inspired by PEPs to evaluate the extent of easy-to-hard generalization achieved by MetaLint.
+    We construct a benchmark of hard-to-detect, broadly relevant best practices inspired by PEPs to probe whether MetaLint models exhibit easy-to-hard generalization.
   3. 3.
-    We benchmark state-of-the-art code and reasoning models on our PEP hard-idiom benchmark and compare them against MetaLint-trained models.
-Our method achieves the highest detection recall and competitive localization scores, even with smaller 4B models and without test-time compute.
+    We benchmark state-of-the-art code and reasoning models on our PEP hard best practice benchmark and compare them against MetaLint-trained models.
+With our training Qwen3-4B outperforms larger models like o3-mini achieving the highest detection recall and comparable localization scores without test-time compute.
   4. 4.
-    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), test-time reasoning settings (with and without CoT), and model scales (3B–8B).
+    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), reasoning settings (with and without CoT), and model scales (3B–8B).
+Overall, MetaLint is not a replacement for linters, but an empirical demonstration that organizing supervision around semantic units enables LLMs to transfer learned code quality reasoning to harder or novel best practices.
+Finally, code quality analysis offers a realistic testbed for studying generalization: while correctness is largely objective, the difficulty progression spans from cases verifiable by automated linters to harder, context-dependent cases requiring human judgment (PEP benchmark).
+This work takes a first step toward such settings, where supervision becomes increasingly incomplete along the difficulty progression.
 2 Related Work
 Code Quality Analysis with Large Language Models.
-A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis.
-Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as rule-guided linters via prompting or fine-tuning.
+A large body of prior work has explored using LLMs for code quality analysis by interfacing them with static analysis tools or fixed rule sets.
+Travis Fischer (2024); lin (2023) treat LLMs as rule-guided linters via prompting or fine-tuning.
 While Blyth et al.
@@
 (2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives.
-LintLLM (Fang et al., 2025) and (Shin et al., 2025) leverages LLMs for linting of Verilog and Quantum computing code.
-Khare et al.
-(2023) show LLMs outperform traditional static analysis tools for security-related CWEs with step-by-step reasoning.
+Fang et al.
+(2025); Shin et al.
+(2025); Khare et al.
+(2023) leverage LLMs for linting and show they can outperform traditional static analysis tools.
 Vijayvergiya et al.
 (2024) train LLMs for best practice violation detection and localization, while Rasheed et al.
-(2024) design a multi-agent review pipeline for maintainability, efficiency, and bugs.
+(2024) design a multi-agent review pipeline for code review.
 Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation.
 Naik et al.
-(2024) and Jaoua et al.
-(2025) integrate LLMs with linters to produce more informative code reviews.
-RIdiom (Zhang et al., 2024c) introduces a rule-based way to identify and refactor non-idiomatic Python code with AST rewrite rules, while Zhang et al.
-(2024b) combines LLMs and rule-based detectors but doesn't explore nuanced idioms like PEP 506 or training LLMs to keep up with evolving best practices.
-Finally, CoUpJava (Jiang et al., 2025a) presents Java version upgrade benchmarks, conceptually similar to our hard PEP idiom benchmark for Python.
-Although prior work demonstrates the potential of LLMs for code quality tasks, it focuses on fixed rule sets or best practices that require retraining as they evolve.
-In contrast, we train models to interpret high-level specifications and perform static analysis, enabling broader generalization.
-Instruction Following for Generalization.
-Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning fixed tasks.
-Prior work shows diverse task instructions allow models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022).
-Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020).
-Instructions serve as high-density task representations, substituting supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022).
-Studies also show that instruction diversity drives generalization, with varied instructions outperforming repeated exposure to identical formats (Charton et al., 2024).
-This phenomenon holds across domains, including program synthesis where task-level prompting facilitates generalization in code generation models (Niu et al., 2023).
-SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness, but relying entirely on LLM-generated data without verifiers.
-These results suggest instruction tuning acts as task-level meta-learning, enabling models to adapt to new tasks through natural language.
-Building on this we model specific code quality idioms as individual tasks and generate large-scale synthetic data for each meta-task to support cross-idiom generalization.
-This allows the trained model to keep pace with new idioms and evolving best practices.
-We also discuss additional related work on easy-to-hard generalization in Appendix B.
+(2024); Kapadnis et al.
+(2025); Jaoua et al.
+(2025) provide linter results to LLMs for more informative code reviews.
+Zhang et al.
+(2024c, b) explore AST rewrite rules and hybrid approaches combining LLMs and rules for code quality refactoring.
+Finally, CoUpJava (Jiang et al., 2025a) propose a Java version upgrade benchmark, mined using AST refactoring rules.
+Across these approaches, static analysis tools and handcrafted rules typically act as either the final decision mechanism or the primary training signal, anchoring model behavior to a fixed set of best practices.
+In contrast, we use static analysis as supervision for the easy end of a difficulty progression, and train models to detect best practice violations from high-level descriptions, enabling transfer to harder and previously unseen cases that extend beyond linter coverage.
+Easy-to-Hard Generalization.
+Recent work suggests that easy-to-hard generalization arises not merely from curriculum ordering, but from how supervision is structured to promote abstraction across related tasks (Chang & Bisk, 2025).
+In mathematical reasoning, models trained on easier problems (e.g., levels 1–3) consistently generalize better to harder benchmarks (e.g., levels 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025), with several studies emphasizing the role of high-quality supervision, particularly for harder instances (He et al., 2024).
+Sun et al.
+(2024) show that reward models trained on simple code and math problems improve performance on more complex tasks.
+More broadly, (Hu et al., 2025; Gaunt et al., 2016) demonstrate that training on structurally simpler instances can yield robustness to longer or more complex reasoning problems, including code.
+Similarly, Zhang et al.
+(2024a) evaluate reward models on algorithmic tasks such as string manipulation and observe transfer from simpler to more complex formats.
+Most closely, Chang & Bisk (2025) argue that easy-to-hard generalization is primarily driven by the learning paradigm and finding the right hypothesis class, rather than data scarcity, model scale or inference limits.
+While prior work studies this phenomenon in domains such as counting, mathematics, or contest coding, code quality analysis offers a complementary setting where correctness is largely objective but supervision becomes increasingly incomplete along the difficulty progression.
+Additional related work on leveraging instruction following for generalization is discussed in Appendix D.
 3 Method
-We design the MetaLint framework to teach an LLM to operationalize idiom descriptions provided in context, rather than memorizing specific idioms, thereby enabling adaptation to novel idioms at test time.
-We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$.
+We design the MetaLint framework to reorganize supervision for best practice violation detection, shifting the hypothesis class from memorizing fixed rules to identifying violations from high-level natural language descriptions of problematic code idioms, in line with recent work showing that such reorganization is critical for easy-to-hard generalization (Chang & Bisk, 2025).
+We formulate best practice violation detection as an instruction-following meta-task $M_{I}$, where for a given problematic idiom $I$, the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$.
 The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$.
-This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$.
-By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices.
+This setup discourages rote memorization by construction, as correctness is defined relative to the provided specification, and encourages adaptive reasoning over the prompt’s description.
+For example flagging best practice violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$.
+By framing best practices as meta-tasks, this approach supports the LLM in remaining flexible and better aligned with evolving best practices.
 We describe the components of our training framework in Figure 1 and Figure 2 below.
 3.1 Synthetic Data Generation
-One of the main goals of our meta-task formulation is enabling easy-to-hard generalization.
-We train LLMs on a set of ``easy'' idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter).
-Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand and detect code idioms from in-context descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to the untrained model.
+One of the main goals of our meta-task formulation is to change the learning task to detecting best practice violatons from high level specifications to encourage transfer from easy to hard cases.
+We train LLMs on a set of ``easy'' best practice idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter).
+Our hypothesis is that training on $I_{\mathcal{L}}$ exposes the model to a structured family of best practice idiom specifications, which supports inductive transfer to hard to detect best practice idioms in $I_{\mathcal{L^{\prime}}}$.
 Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision.
@@
 Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples.
-An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix C.1.
+An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix E.1.
 For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.
 3.2 Instruction Supervised Fine-Tuning
-As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$.
+Instruction fine-tuning serves to align the model's outputs with the natural language best practice specifications using the linter outputs as a reference for easy cases.
+We train the target LLM $\Phi$ on a set of linter-detectable, easy best practice idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$.
 The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$.
-The model's output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix C.1).
+The model's output is a list of best practice violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix E.1).
 In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND.
-We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python data, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data.
+This formulation explicitly penalizes hallucinated violations of best practices not in the prompt, encouraging the model to flag only violations of the provided best practice and discouraging memorization.
+We try to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data.
 This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
 3.3 Verifiable Reward Model and Preference Optimization
-For preference optimization, we adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model.
+We use preference optimization to further improve consistent application of the best practice specification by contrasting accurately localized violations with inaccurate ones, as determined by linter results.
+We adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model.
 It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2).
@@
 Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap.
-Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
+Since each meta-task $M_{I}$ corresponds to a single best practice idiom $I$, we compute one F1-score (reward) per instance.
 Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity.
@@
 Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$.
-For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$.
+For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{w}=y_{i}$, $y_{l}=y_{j}$, and store the instance $(x,y_{w},y_{l})\in\mathcal{D}_{p}$.
+By constructing contrastive pairs within the same meta-task, RS-DPO encourages the model to apply a given specification consistently across variations in code structure, rather than across best practices.
 Following Khaki et al.
 (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
-$\Phi^{RL}=\arg\max\sum_{(x,y_{win},y_{lose})\in\mathcal{D}_{p}}\log\sigma\left(\beta\log\frac{\Phi^{RL}(y_{win}|x)}{\Phi^{SFT}(y_{win}|x)}-\beta\log\frac{\Phi^{RL}(y_{lose}|x)}{\Phi^{SFT}(y_{lose}|x)}\right)$
+$\Phi^{RL}=\arg\max\sum_{(x,y_{w},y_{l})\in\mathcal{D}_{p}}\log\sigma\Big($ $\beta\log\frac{\Phi^{RL}(y_{w}\mid x)}{\Phi^{SFT}(y_{w}\mid x)}$
+$-\beta\log\frac{\Phi^{RL}(y_{l}\mid x)}{\Phi^{SFT}(y_{l}\mid x)}\Big)$
 Here, $\sigma$ denotes the sigmoid function, and $\beta=0.1$ is the KL penalty coefficient, corresponding to low-to-moderate regularization.
 3.4 Training with Reasoning Traces
-Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning.
+Reasoning traces can act as explicit intermediate representations to support transfer across best practice idioms.
 To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection.
@@
 4.1 Evaluation Metrics
-We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs.
+We evaluate the LLM's ability to detect best practice violations through two tasks: detection, which assesses whether a best practice is violated at least once in a code file, and localization, which evaluates whether the model accurately identifies the lines with the problematic idiom.
 For both tasks, we report precision, recall, and F-score metrics.
-Detection metrics are calculated at the corpus level for each idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers.
+Detection metrics are calculated at the corpus level for each best practice idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers.
 To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics).
-For localization, metrics are averaged only across instances with at least one line of idiom violation in the ground truth.
-Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix D.1.
-4.2 Generalization on Synthetic Data
-To evaluate whether MetaLint training produces adaptive LLMs that handle evolving best practices and novel idioms at test time, we explore transfer settings spanning Python & Java.
+For localization, metrics are averaged only across instances with at least one violation in the ground truth.
+Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix F.1.
+4.2 Generalization across Easy Best Practices
+To evaluate whether MetaLint training encourages transfer to novel easy-to-detect best practices at test time, we explore transfer settings spanning Python & Java.
 Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1.
 The data has a 74:26 no-violation-to-violation split, similar to the SFT training set.
-Idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
-In domain. 5 idioms identical to those in SFT training, to assess whether MetaLint improves performance on explicitly trained idioms.
+Best practice idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
+In domain. 5 idioms seen during training, to assess whether MetaLint improves performance on in-domain idioms.
 Near transfer. 10 idioms with specifications similar but not identical to training idioms to probe memorization. Reliance on memorized patterns, may hurt performance due to interference.
@@
 PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries.
-We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer between PMD and JEP idioms to test adaptation to novel Java idioms.
+We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer from PMD to JEP and vice-versa.
 These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
-4.3 PEP Hard Idiom Benchmark
-Benchmark Construction: To test whether MetaLint helps LLMs interpret high-level idiom specifications and generalize to nuanced idioms that linters miss, we construct a benchmark of ``hard idioms'' from 15 PEPs defining semantic or abstract behaviors beyond syntax.
-We design heuristics per PEP (Table 13, 14 and 15) to detect guideline violations and search the STACK-V2 corpus, prioritizing recall to retrieve broad candidate sets for manual selection.
-These idioms cannot be reliably detected by simple pattern matching, making them ideal for evaluating model's true understanding versus rote memorization.
-From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (“start” and “end”) of the violated code, providing ground-truth for localization.
+4.3 PEP Hard Best Practice Benchmark
+To test whether MetaLint trained LLMs exhibit easy-to-hard generalization at test time, we construct a benchmark of hard-to-detect best practices from 15 PEPs that require abstract, context-dependent reasoning.
+We design heuristics per PEP (Table 15, 16 and 17) to find candidate violations in the STACK-V2 corpus, prioritizing recall to retrieve a broad set for manual selection.
+These best practices violations cannot be reliably detected by rule-based pattern matching, making them ideal for evaluating a model's contextual reasoning capabilities.
+From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (``start'' and ``end'') of the violated code, providing ground-truth for localization.
 We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases.
-The final benchmark contains 536 examples (52% violations, 48% violation-free), enabling evaluation of MetaLint’s generalization from easy to hard Python idioms.
-Evaluating Easy-to-Hard Generalization: We use the PEP hard idiom benchmark to test whether training on synthetic data for linter-detectable idioms improves performance on hard idioms.
-We evaluate the base model, SFT, and DPO-trained models on this benchmark.
-Benchmarking on Hard Idioms: We evaluate state-of-the-art open and closed-source code and reasoning LLMs on the PEP hard idiom benchmark, comparing them to MetaLint-trained models.
-Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B Agarwal et al.
-(2025).
-Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 OpenAI (2025b).
+The final benchmark contains 536 examples (52% violation, 48% no-violation), enabling evaluation of easy-to-hard generalization by comparing the base model, against the SFT, and DPO-trained models.
+We also compare MetaLint models trained on Ruff best practices against state-of-the-art open and closed-source code and reasoning LLMs.
+Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B (Agarwal et al., 2025).
+Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 (OpenAI, 2025b).
 We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
 5 Results
-To test whether MetaLint training leads to cross-idiom generalization instead of mere memorization of the training idioms and whether it can produce models that can keep up with evolving code quality standards, we present the transfer performance on the synthetic data for ``easy'' idioms in section 5.1.
-Then we explore the extent to which MetaLint training achieves easy-to-hard generalization from the synthetic easy idioms to hard, manually curated PEP idioms in section 5.2.
-Finally, we compare MetaLint trained models against state-of-the-art code and reasoning models on the manually curated hard PEP idioms in section 5.3.
-5.1 Generalization on Synthetic Data
+We evaluate the effectiveness of MetaLint along three dimensions: (1) generalization to unseen easy best practices (Section 5.1); (2) easy-to-hard generalization across best practices (Section 5.2); and (3) whether MetaLint enables a 4B model to achieve performance competitive with state-of-the-art code and reasoning models (Section 5.3).
+5.1 Generalization across Easy Best Practices
 Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 21).
 While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 18.
-The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop in the case of Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics at the cost of a slight drop in detection precision.
-We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix D.3.
-Additionally, Table 16 shows that while SFT can lead to slight gains for the transfer settings (near transfer and far transfer), most gains emerge in the DPO stage, especially for non-reasoning models and detection recall.
-Overall this suggests that SFT can lead to memorization of the training idioms while DPO leads to generalization to novel idioms.
+The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop for Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics with a slight drop in detection precision.
+We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix F.3.
+Additionally, Table 18 shows that while SFT can lead to slight gains for the transfer settings, most gains emerge in the DPO stage, especially for non-reasoning models and detection recall.
+Overall this suggests that SFT is prone to memorizing best practices used for training, while DPO generalizes to novel best practices.
+PMD and JEP Tree-Sitter Idioms: To evaluate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 31).
+Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO.
+For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct.
+Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP.
+In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores, since our metrics exclude that class for detection and only score positive cases for localization.
+Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5).
+In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD's broader diversity and more challenging specifications (Appendix E.5).
+Overall, MetaLint outperforms the base model on novel best practice violations, though gains depend on the diversity of training idioms and the gap in instruction quality between training and test data.
+Table 1: Cross-Idiom Generalization on Python Ruff Idioms: Effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning).
+Best score across the compared training setups per model are bolded.
 Model Detection Localization
@@
 Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.7710 0.7571 0.7640
-Table 1: Cross-Idiom Generalization on Python Ruff Idioms: Effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning).
+Table 2: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 31).
+The transfer column indicates training and test data on the left and right side of the arrow.
 Best score across the compared training setups per model are bolded.
-PMD and JEP Tree-Sitter Idioms: To demonstrate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 28).
-Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO.
-For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct.
-Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP.
-In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores because our metrics exclude that class for detection and only score positive cases for localization.
-Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5).
-In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD’s broader diversity and more challenging specifications (Appendix C.5).
-Overall, MetaLint training consistently yields more adaptable models than the base model, but performance depends on the diversity of training idioms and the gap in instruction quality between training and test data.
 Model Transfer Detection Localization
@@
 Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
-Table 2: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28).
-The transfer column indicates training and test data on the left and right side of the arrow.
+5.2 Evaluating Easy-to-Hard Generalization
+To evaluate whether MetaLint training on easy to detect Ruff best practices improves performance on hard, PEP best practices, we report results on our PEP hard best practice benchmark (Table 3, full results in Table 22).
+At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT is prone to memorization of the training distribution.
+In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix G.2).
+An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited.
+Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to encourage easy-to-hard generalization.
+Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision.
+Our analysis attributes the CoT model's reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix G.3.
+Finally we analyze the impact of the illustrative examples in the prompt in Table 19.
+Table 3: Easy-to-Hard Generalization: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3).
 Best score across the compared training setups per model are bolded.
-5.2 Evaluating Easy-to-Hard Generalizations
-To evaluate whether MetaLint training on easy, linter-detectable Ruff idioms improves performance on hard, manually curated PEP idioms, we report results on our PEP hard idiom benchmark (Table 3, full results in Table 19).
-At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT can induce memorization of the training distribution and reduce adaptability.
-In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix E.2).
-An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited.
-Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to unlock easy-to-hard generalization.
-Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision.
-Our analysis attributes the CoT model’s reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix E.3.
 Model Detection Localization
@@
 Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
-Table 3: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
-Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3).
-Best score across the compared training setups per model are bolded.
-5.3 Benchmarking on Hard Idioms
-Table 4 compares the best-performing Qwen3-4B MetaLint DPO models against state-of-the-art code and reasoning models (full results in Table 17).
-Detection: In terms of detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini).
-However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
-Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix E.2) and outperform GPT-oss-20B.
-This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy idioms, and that the non-CoT model does not use test-time compute.
-Overall, the strong results, especially the best-in-class recall of the non-CoT model, demonstrate the effectiveness of our framework in achieving easy-to-hard generalization.
-This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction fine-tuning and DPO rather than relying on rote memorization.
+Table 4: Benchmarking on Hard Best Practices: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training.
+The best scores are bolded and second best and underlined.
 Detection Localization
@@
 GPT-5 (high) 0.9130 0.5673 0.6998 0.4397 0.4257 0.4326
-Table 4: Benchmarking on Hard Idioms: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training.
-The best scores are bolded and second best and underlined.
+5.3 Benchmarking on Hard Best Practices
+Table 4 compares the best-performing Qwen3-4B MetaLint DPO models against state-of-the-art code and reasoning models (full results in Table 20).
+Detection: For detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini).
+However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
+Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix G.2) and outperform GPT-oss-20B.
+This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy best practices, and that the non-CoT model does not use test-time compute.
+We also identify localization errors and their frequency in Table 32.
+Overall, our models perform comparably to state-of-the-art systems, with the non-CoT Qwen3-4B variant achieving the highest recall, and both MetaLint-trained Qwen and LLaMA models exhibiting evidence of easy-to-hard generalization.
+This aligns with Chang & Bisk (2025), which highlights the role of reorganizing supervision and learning paradigm, rather than scaling or optimization alone, in enabling generalization along difficulty progressions.
 6 Conclusion and Future Work
-Our results show that MetaLint training fosters adaptive reasoning over idiom specifications rather than rote memorization.
-We observe generalization to unseen idioms in Python and Java, across three linters (Ruff, PMD, JEP tree-sitter), two model families (Qwen, Llama), reasoning and non-reasoning settings, and multiple scales (3B, 4B, 8B).
-Easy-to-hard generalization occurs from linter-detectable Ruff idioms to harder PEP idioms, with SFT teaching output formatting and DPO enabling true generalization.
-Compared to state-of-the-art code and reasoning models, MetaLint-trained Qwen models have detection comparable with o3-mini and GPT-5, achieving highest recall (non-CoT) and third-best precision (CoT).
-Localization lags but surpasses GPT-oss-20B with only 4B parameters and no test-time compute and is comparable to o3-mini, demonstrating efficiency.
-These results highlight the effectiveness of instruction fine-tuning and preference optimization on synthetic data for reasoning and generalization, even with scarce annotated examples.
-For mechanically easy idioms, linters remain cost-effective, but MetaLint enables detection of abstract idioms, supporting personalized, evolving code quality standards.
-We plan to release code and data for reproducibility.
-Future work includes training for automated refactoring and exploring advanced RL methods like Group Relative Policy Optimization (GRPO) Shao et al.
-(2024).
-Acknowledgments
+Our results show that MetaLint encourages inductive transfer over best practice specifications, rather than memorization of fixed rules.
+We observe generalization to unseen best practices across Python and Java, multiple linters, model families, reasoning settings, and model scales.
+Models trained with MetaLint also exhibit easy-to-hard generalization, transferring from linter-detectable best practices to harder PEP violations.
+With only 4B parameters, MetaLint-trained Qwen models achieve detection performance comparable to strong code and reasoning systems, achieving the highest recall and competitive precision.
+Localization remains more challenging but is competitive with o3-mini and surpasses GPT-oss-20B without test-time compute.
+Overall, these results support the view that reorganizing supervision, rather than scaling alone, enables inductive generalization beyond the training distribution (Chang & Bisk, 2025).
+MetaLint is not a replacement for static analysis, but a proof of concept that inductive learning can improve LLM performance along a natural difficulty progression in code quality analysis.
+Future work will target automated refactoring and better RL training like Group Relative Policy Optimization (GRPO) (Shao et al., 2024).
+Acknowledgements
 This work was partially supported by a generous collaborative research grant from Oracle Labs.
 We are grateful to Jack Sullivan, Qinlan Shen, Adam Pocock, and Nidhi Chandra at Oracle Labs for their guidance, feedback, and insightful discussions that played a key role in shaping the core ideas behind MetaLint and advancing the project.
-Appendix A Limitations
-Despite the promising results achieved by MetaLint, our work has some limitations that we plan to address in future research.
-For the CoT setting, we didn't explore whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model.
-We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data.
-As a result, we adopted a simpler rejection sampling or RS-SFT strategy.
-Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data.
-Our current experiments also focus on training on one language at a time, such as only Python or Java.
-Future work will explore joint training and extension to more programming languages like JavaScript, Ruby, Go, etc., as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa.
-Finally, while we do not evaluate or train for refactoring of the idiom-violating code, we plan to do so in future work.
-Appendix B More Related Work
-Easy-to-Hard Generalization.
-Research shows that training on simpler problems enhances generalization to harder ones in math, algorithms, and code, motivating its application to code quality analysis.
-In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025).
-Several works emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024).
-Beyond math, Sun et al.
-(2024a) shows that reward models trained on simple code and math problems improve performance on complex ones.
-Broader studies on multi-task and length generalization (Hu et al., 2025) and differentiable programming (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code.
-Zhang et al.
-(2024a) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats.
-Drawing inspiration from this work, we train MetaLint on large-scale synthetic data covering easily detectable code idioms handled by rule-based linters, and hypothesize that these simple patterns serve as stepping stones toward generalizing to complex, novel PEP idioms.
-Appendix C Method Additional Details
-C.1 MetaLint Instruction Following Prompt
-We used the following instruction following style prompt to train the model with synthetic Ruff idiom data for the meta-linting task:
+Impact Statement
+This work aims to improve the reliability and generalization of automated code quality analysis by enabling models to adapt to evolving and previously unseen best practices.
+By reducing reliance on brittle, static rule sets, MetaLint has the potential to lower maintenance costs for software tooling and improve developer productivity and security outcomes, particularly in settings where best practices change rapidly.
+However, as with any automated code analysis system, incorrect or overconfident recommendations could mislead developers or create false assurance if deployed without human oversight.
+We mitigate this risk by grounding supervision in existing static analysis signals and framing outputs as assistive diagnostics rather than authoritative judgments.
+We do not anticipate significant negative societal impacts beyond those common to developer tooling, and we expect the primary effects of this work to be beneficial.
+Appendix A Appendix
+We organize the Appendix in five parts: 1) Limitations of our work (Appendix B), 2) Clarifications and definitions of terms used in the paper (Appendix C), 3) Additional related work about instruction following and generalization (Appendix D), 4) Method details like prompts and hyperparameters (Appendix E), 5) Additiomal experimental details (Appendix F) and 6) Additional results and ablations (Appendix G).
+Appendix B Limitations
+Despite promising results, including evidence of easy-to-hard generalization, MetaLint has several limitations that suggest directions for future work:
+  • Chain-of-Thought (CoT) reasoning: We did not explore whether non-CoT models can be trained to generate CoT-style reasoning using supervision from a teacher model.
+  • Self-improvement and data generation: We attempted self-improvement strategies for RS-SFT data generation (e.g., STaR (Zelikman et al., 2022)) when the base model failed.
+However, generating CoTs that do not directly reference provided hints proved challenging and risks contaminating the training data.
+We therefore adopted a simpler rejection sampling (RS-SFT) strategy.
+  • Preference optimization methods: We only experimented with DPO for preference optimization.
+Alternative approaches, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024), combined with our verifiable linter-based reward model, could be explored.
+  • Difficulty progression and curriculum: Our current setup does not implement a fine-grained difficulty progression or a curriculum over meta-task abstractions.
+Incorporating such structure could encourage more systematic induction over code idioms and code quality concepts.
+  • Multilingual training: Experiments focus on one programming language at a time (Python or Java).
+Future work could explore joint training across multiple languages, cross-language transfer, and structured learning that encourages abstraction across similar language constructs (e.g., Python’s PEP 557: Data Classes vs.
+Java’s JEP 395: Records).
+  • Task scope: We focus exclusively on best practice violation detection and do not yet explore inductive transfer over code refactoring or other code quality tasks.
+Appendix C Clarification and Term Definitions
+Because this paper studies a software engineering–heavy problem setting, we clarify the key technical terms used throughout.
+Code Idiom / Programming Idiom.
+A code idiom is a recurring, language-specific pattern of code that conveys a particular semantic intent or programming practice beyond what is expressed by its literal syntax.
+Code Quality Analysis.
+Code quality analysis refers to the systematic evaluation of source code with respect to software engineering attributes such as readability, complexity, maintainability, and adherence to best practices, using automated tools or model-based methods.
+Although best practice analysis is only one component of code quality analysis, we use these terms interchangeably in this paper for brevity.
+Best Practice Violation Detection.
+Best practice violation detection is a subproblem of code quality analysis that involves identifying and localizing code fragments that violate established or recommended programming practices.
+In this work, we operationalize best practice violation detection as the task of identifying problematic code idioms, i.e., recurring patterns that correspond to non-recommended practices.
+While this formulation is more restrictive than the full space of best practices, it enables a systematic task structure that supports controlled supervision and generalization.
+Additionally we often use the shorthand best practice idiom or idiom to refer to the problematic idiom that needs to be detected to flag a best practice violation.
+Appendix D More Related Work
+Instruction Following and Generalization.
+Instruction tuning has emerged as a powerful form of meta-learning, enabling cross-task generalization by training models to interpret and follow natural language instructions rather than memorizing fixed tasks.
+Prior work shows that exposure to diverse task instructions allows models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022).
+Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020).
+Instructions serve as high-density task representations, substituting for explicit supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022).
+Studies also show that instruction diversity drives generalization: varied instructions outperform repeated exposure to identical formats (Charton et al., 2024).
+This effect holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023).
+SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness but relying entirely on LLM-generated data without verifiers.
+These results suggest that instruction tuning functions as task-level meta-learning, enabling models to adapt to new tasks through natural language.
+Building on this insight, we model individual code quality idioms as separate meta-tasks and generate large-scale synthetic data for each task to support cross-idiom generalization.
+This approach allows the trained model to adapt to new idioms and evolving best practices.
+Appendix E Method Additional Details
+E.1 MetaLint Instruction Following Prompt
+We used the following instruction following style prompt to train the model with synthetic Ruff best practice idiom data for the meta-linting task:
 MetaLint Instruction Following Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter.
@@
 Code file: {CODE_FILE} Violations per idiom:
-An example input with the code file and idiom spec populated as well as the expected JSON style output is shown below:
+An example input with the code file and best practice idiom spec populated as well as the expected JSON style output is shown below:
 Example Ruff Meta-Task Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations.
@@
 event, *args, **kwargs):", "fix": null}
-C.2 DPO Contrastive Pair and RS-SFT Sampling Details
+E.2 DPO Contrastive Pair and RS-SFT Sampling Details
 To generate RS-DPO contrastive samples (or RS-SFT outputs) from the baseline SFT (or untrained) models, we used the following hyperparameters: nucleus sampling with a maximum of 2048 new tokens, $k=5$ sampled outputs per input, temperatures picked cyclically from {0, 0.3, 0.5, 0.7, 1}, a top-$p$ (cumulative probability threshold) of 0.95, and a seed of $42+i$, where $i\in\{1,\dots,k\}$, to encourage both reproducibility and output diversity.
 For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models.
-In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section C.4.
-C.3 Training Hyperparameters and Computational Environment
+In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section E.4.
+E.3 Training Hyperparameters and Computational Environment
 Python SFT/RS-SFT hyperparameters:
@@
 Inference used a similar setup with GPU allocation varying by model size.
-C.4 Baseline Inference Details
+E.4 Baseline Inference Details
 We use the following hyperparameters for performing inference with the baseline LLMs:
@@
 Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches.
-For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Idiom Benchmark.
+For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Best Practice Benchmark.
 Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter.
@@
 and as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:** NO VIOLATIONS FOUND Violations per idiom:
+Table 5: JEP Best Practice Idiom Specifications (1/3): This table presents 15 best practice idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns.
+The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern.
+The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag problematic idioms for synthetic data creation.
 JEP# JEP Title Definition Example(s) Tree Sitter Queries
@@
 jep_361_before_custom_detectors
-Table 5: JEP Idiom Specifications (1/3): This table presents 15 idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns.
-The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern.
-The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag idioms for synthetic data creation.
+Table 6: JEP Best Practice Idiom Specifications (2/3)
 JEP# JEP Title Definition Example(s) Tree Sitter Queries
@@
 ) @jep_395_after_record_…
-Table 6: JEP Idiom Specifications (2/3)
+Table 7: JEP Best Practice Idiom Specifications (3/3)
 JEP# JEP Title Definition Example(s) Tree Sitter Queries
@@
 )) @jep_323_after_local_var_lambda
-Table 7: JEP Idiom Specifications (3/3)
-C.5 PMD Idiom Specifications
+E.5 PMD Best Practice Idiom Specifications
 We scrape PMD idioms specification from the Java section of the PMD rules documentationhttps://docs.pmd-code.org/latest/pmd_rules_java.html.
-The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as idiom violations, as can be seen in the example below.
+The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as best practice idiom violations, as can be seen in the example below.
 PMD Rule Specification: UnitTestShouldIncludeAssert
@@
 </rule>
+Table 8: List of JEPs addressed by our tree-sitter synthetic data.
+The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features.
+The Before and After columns indicate whether we include rules/patterns to flag the old problematic idiom or new recommended idiom introduced by the JEP.
 JEP # Before After Title JDK# Release Date
@@
 10 20 Mar 2018
-Table 8: List of JEPs addressed by our tree-sitter synthetic data.
-The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features.
-The Before and After columns indicate whether we include rules/patterns to flag the old idiom or new idiom introduced by the JEP.
 Java MetaLint Instruction Following Prompt Task Instructions (1/2): Look at the following code idiom specification with definitions and examples: {IDIOM_SPEC} Task Instructions (2/2): Given this idiom, your task is to look at a code file and detect violations of the above idiom, and flag them like a linter.
@@
 Code file: {CODE_FILE} Violations per idiom:
-Appendix D Additional Experimental Details
-D.1 Evaluation Metrics
-Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations.
+Appendix F Additional Experimental Details
+F.1 Evaluation Metrics
+Let $I$ denote an best practice idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations.
 For each dataset instance with input prompt $x$ and ground truth set of line numbers $y$, $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$.
@@
 $P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
-D.2 Idioms Chosen for Ruff Idiom Transfer Dataset
-Table 9 lists the Ruff idioms used in the SFT training and synthetic transfer evaluation test sets.
+F.2 Idioms Chosen for Ruff Best Practice Transfer Dataset
+Table 9 lists the Ruff best practice idioms used in the SFT training and synthetic transfer evaluation test sets.
 Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
+Table 9: Ruff best practice idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
 Training Set Idioms Test Set Idioms
@@
 S201, S202, S301, S302, S303
-Table 9: Ruff idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
 Figure 3: ID: In-Domain, NeT: Near Transfer, FaT: Far Transfer.
-D.3 DPO No Violation Fraction Ablations
+F.3 DPO No Violation Fraction Ablations
 We analyze the impact of varying the amount of samples with zero violations used for RS-DPO training.
@@
 Based on these insights, we conducted a limited ablation on the CoT model, evaluating 2% and 5% inclusion to determine the optimal setting for both detection and localization (as shown in Table 11).
-Fraction of
-NV data
-Detection Localization
+Table 10: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT.
+Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND.
+Conversely, including 100% improves precision but leads to reduced recall and localization performance.
+All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Best Practice Transfer test set.
+Fraction of NV data Detection Localization
 $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
@@
 100% 0.8565 0.4152 0.5593 0.4041 0.4056 0.4048
-Table 10: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT.
-Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND.
-Conversely, including 100% improves precision but leads to reduced recall and localization performance.
-All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Idiom Transfer test set.
-Fraction of
-NV data
-Detection Localization
+Table 11: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model with CoT.
+We perform limited ablations because of the insights from the non CoT model training.
+Fraction of NV data Detection Localization
 $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
@@
 5% 0.9234 0.8643 0.8929 0.771 0.7571 0.764
-Table 11: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model with CoT.
+Table 12: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Llama3.2-3B-Instruct model.
 We perform limited ablations because of the insights from the non CoT model training.
-Fraction of
-NV data
-Detection Localization
+Fraction of NV data Detection Localization
 $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
@@
 2% 0.6636 0.6057 0.6333 0.4869 0.4745 0.4806
-Table 12: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Llama3.2-3B-Instruct model.
-We perform limited ablations because of the insights from the non CoT model training.
-D.4 PEP Benchmark Creation Additional Details
-As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP idioms. These are summarized in Table 13, 14 and 15.
+Table 13: Effect of Simple Ensembling Strategies on Detection and Localization Performance: Comparison of intersection and union-based ensembling of SFT and RS-SFT models on Qwen3-4B.
+The intersection strategy retains only line numbers predicted by both models, while the union strategy aggregates all predicted line numbers from either model.
+Both ensembling approaches underperform the individual models, with the intersection strategy exhibiting particularly poor precision and recall, suggesting that the observed precision–recall differences are driven by training setup rather than a fundamental trade-off.
+Ensembling Strategy Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B DPO Intersection (non-CoT & CoT) 0.0667 0.0033 0.0063 0.0000 0.0000 0.0000
+Qwen3-4B DPO Union (non-CoT & CoT) 0.6188 0.4187 0.4995 0.1952 0.1251 0.1524
+Table 14: Effect of Removing NO VIOLATIONS (NV) Data on Hard PEP Benchmark Performance: Full detection and localization results for Qwen3-4B trained with SFT+RS-DPO under different NV data settings.
+The 0% NV variant achieves substantially higher recall than the model reported in the main paper, with only a moderate reduction in precision.
+Since the precision drop is smaller than the recall gain, this leads to an improved overall F score, confirming that the originally reported recall was an underestimate and further strengthening our state-of-the-art recall claim.
+Model Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B (SFT + RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
+Qwen3-4B (SFT + RS-DPO, 0% NV) 0.6658 0.8037 0.7283 0.3797 0.2063 0.2674
+Qwen3-4B w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+F.4 PEP Benchmark Creation Additional Details
+As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP best practice idioms. These are summarized in Table 15, 16 and 17.
+Table 15: High recall heuristics used to find instances of PEP violations that human annotators vet
 PEP Description Heuristics Example
@@
 process(match)
-Table 15: High recall heuristics used to find instances of PEP violations that human annotators vet
-Appendix E More Results
+Appendix G More Results
+Table 18: Cross-Idiom Generalization on Python Ruff Best Practice Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff best practice idioms, and the performance is reported on other Ruff best practice idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
 Model In-Domain Near Transfer Far Transfer
@@
 +RS-DPO 0.75 (+0.21) 0.92 (+0.49) 0.83 (+0.35) 0.81 (+0.12) 0.71 (+0.03) 0.76 (+0.07) 0.61 (+0.14) 0.59 (+0.08) 0.60 (+0.11)
-Table 16: Cross-Idiom Generalization on Python Ruff Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
-Models are trained on easy synthetic Python Ruff idioms, and the performance is reported on other Ruff idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
-E.1 Expanded Results on the PEP Hard Idiom Benchmark
+Table 19: Ablation on In-Context Examples for the Hard PEP Benchmark: Effect of including before/after code examples in the prompt for base Qwen3-4B models and final MetaLint models trained with SFT and DPO.
+Across most settings, removing examples results in minimal changes in detection and localization performance, indicating that the natural-language descriptions alone are often sufficient.
+The only notable degradation occurs for the CoT-trained MetaLint model, which exhibits a modest drop in detection performance without examples due to an increased tendency to predict NO VIOLATIONS, suggesting that this more conservative variant benefits more from additional in-context guidance.
+Model Examples? Detection Localization
+$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
+Qwen3-4B Yes 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
+Qwen3-4B w CoT Yes 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
+Qwen3-4B (SFT + RS-DPO) Yes 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
+Qwen3-4B w CoT (RS-SFT + RS-DPO) Yes 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
+Qwen3-4B No 0.5556 0.1736 0.2645 0.0802 0.0733 0.0766
+Qwen3-4B w CoT No 0.8304 0.3669 0.5089 0.2572 0.1542 0.1928
+Qwen3-4B (SFT + RS-DPO) No 0.7167 0.7268 0.7217 0.3396 0.1909 0.2444
+Qwen3-4B w CoT (RS-SFT + RS-DPO) No 0.8215 0.4249 0.5601 0.2860 0.1983 0.2342
+G.1 Expanded Results on the PEP Hard Best Practice Benchmark
 We show the expanded results across various model sizes for the evaluated model families in Table 20.
@@
 We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models.
-This suggests that SFT alone may lead to overfitting on the Ruff idiom distribution and struggles to generalize from easy to hard cases without DPO training.
+This suggests that SFT alone may lead to overfitting on the Ruff best practice idiom distribution and struggles to generalize from easy to hard cases without DPO training.
 These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline.
@@
 This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
+Table 20: Results on the hard PEP benchmark to measure easy to hard generalization.
 Detection Localization
@@
 GPT-5 (high) 0.913 0.5673 0.6998 0.4397 0.4257 0.4326
-Table 17: Results on the hard PEP benchmark to measure easy to hard generalization.
+Table 21: Cross-Idiom Generalization on Python Ruff Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff best practice idioms and tested on other Ruff best practice idioms with varying levels of transfer (section 4.2).
+Best score across the compared training setups per model is bolded.
 Model Detection Localization
@@
 Llama3.2-3B-Instruct + SFT + RS-DPO 0.654 0.6468 0.6504 0.491 0.4788 0.4848
-Table 18: Cross-Idiom Generalization on Python Ruff Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
-Models are trained on easy synthetic Python Ruff idioms and tested on other Ruff idioms with varying levels of transfer (section 4.2).
+Table 22: Easy-to-Hard Generalization on PEP Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
+Models are trained on easy synthetic Python Ruff best practice idioms and tested on hard manually curated PEP best practice violation detection data which can't be handled by linters or static analyzers (section 4.3).
 Best score across the compared training setups per model are bolded.
@@
 Llama3.2-3B-Instruct + SFT + RS-DPO 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
-Table 19: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B.
-Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3).
-Best score across the compared training setups per model are bolded.
-E.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
+G.2 Statistical Significance of Results on the PEP Hard Best Practice Benchmark
 To analyze the statistical significance of performance differences over the PEP benchmark, we conduct Wilcoxon signed-rank tests comparing various MetaLint variants against each other and against baseline models.
@@
 Also the MetaLint non CoT (SFT+RS-DPO) variant has no singificant difference in performance compared to o3-mini, soldifying that MetaLint without CoT has generalized to the point of being as capable as o3-mini (even though the Qwen3-4B models without CoT and Qwen3-4B model with CoT perform worse than it with the difference being statistically singificant in Table 23).
+Table 23: Wilcoxon signed-rank test results comparing untrained Qwen3-4B variants with o3-mini, using Bonferroni-adjusted significance threshold $\alpha=0.025$.
+Each cell reports the test statistic (p-value).
 Model Comparison Detection Localization P Localization R
@@
 Qwen3-4B w CoT vs o3-mini 921.5 (3.23e-09) 2385.5 (9.61e-02) 1891.0 (4.99e-04)
-Table 20: Wilcoxon signed-rank test results comparing untrained Qwen3-4B variants with o3-mini, using Bonferroni-adjusted significance threshold $\alpha=0.025$.
-Each cell reports the test statistic (p-value).
 Table 27 shows the effect of using a CoT for the Qwen3 model families and we notice that using a CoT leads to singificant gains for all metrics for the 4B and 8B models indicating that for smaller models CoTs might be essential for good performance on this task.
@@
 We only compared GPT-4o and its successor GPT-4.1 and o3-mini against o4-mini and the results show that GPT-4.1 is only significantly better for localization recall while o4-mini is beter than o3-mini for overall localization but not for detection.
+Table 24: Wilcoxon signed-rank test results comparing MetaLint variants.
+Each cell reports test statistic (p-value).
+All the MetaLint models are trained Qwen3-4B variants.
+We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
 Model Comparison Detection Localization P Localization R
@@
 MetaLint (SFT+RS-DPO) vs MetaLint w CoT (RS-SFT+RS-DPO) 8140.0 (5.55e-01) 2568.5 (8.42e-01) 2544.0 (4.44e-01)
-Table 21: Wilcoxon signed-rank test results comparing MetaLint variants.
-Each cell reports test statistic (p-value).
-All the MetaLint models are trained Qwen3-4B variants.
-We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
+Table 25: Wilcoxon signed-rank test results comparing MetaLint models against their untrained counterparts, with Bonferroni-adjusted significance threshold $\alpha=0.0125$.
+Each cell reports the test statistic (p-value).
 Model Comparison Detection Localization P Localization R
@@
 Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) 1728.0 (1.83e-02) 1011.0 (1.53e-03) 966.0 (1.02e-03)
-Table 22: Wilcoxon signed-rank test results comparing MetaLint models against their untrained counterparts, with Bonferroni-adjusted significance threshold $\alpha=0.0125$.
-Each cell reports the test statistic (p-value).
+Table 26: Wilcoxon signed-rank test statistics and p-values comparing MetaLint variants against baseline models.
+All the MetaLint variants are Qwen3-4B variants and Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints.
+We use the Bonferroni corrected significance threshold $\alpha=0.0017$.
 Model Comparison Detection Localization P Localization R
@@
 o3-mini vs MetaLint w CoT (RS-SFT+RS-DPO) 1944.0 (7.15e-04) 3169.0 (5.16e-01) 2683.0 (4.23e-01)
-Table 23: Wilcoxon signed-rank test statistics and p-values comparing MetaLint variants against baseline models.
-All the MetaLint variants are Qwen3-4B variants and Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints.
-We use the Bonferroni corrected significance threshold $\alpha=0.0017$.
+Table 27: Wilcoxon signed-rank test results measuring the effect of Chain-of-Thought (CoT) prompting across Qwen3 model scales.
+Each cell reports test statistic (p-value).
+We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
 Model Comparison Detection Localization P Localization R
@@
 Qwen3-32B vs Qwen3-32B w CoT 1572.5 (2.75e-01) 1767.0 (6.35e-06) 2596.5 (7.31e-02)
-Table 24: Wilcoxon signed-rank test results measuring the effect of Chain-of-Thought (CoT) prompting across Qwen3 model scales.
-Each cell reports test statistic (p-value).
-We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
+Table 28: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings.
+Each cell shows the test statistic (p-value).
+All Qwen2.5 and Qwen2.5Coder variants are instruction-tuned checkpoints.
+We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
 Model Comparison Detection Localization P Localization R
@@
 Qwen2.5-14B vs Qwen2.5-32B 1925.0 (1.87e-04) 1858.5 (6.60e-03) 2108.5 (1.43e-02)
-Table 25: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings.
-Each cell shows the test statistic (p-value).
-All Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints.
-We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
+Table 29: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
 Model Comparison Detection Localization P Localization R
@@
 o3-mini vs o4-mini 575.0 (6.68e-01) 1171.0 (1.44e-03) 1079.0 (1.47e-04)
-Table 26: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
-E.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
+G.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
 We observe that a significant portion of the lower detection recall of the CoT MetaLint Qwen3-4B model, relative to its non CoT counterpart, can be attributed to its higher tendency to predict NO VIOLATIONS FOUND in cases that do, in fact, contain violations.
@@
 While errors span a long tail across many PEPs, the majority are concentrated in three: PEP614, PEP593, and PEP616, which motivates our focused analysis on these cases.
+Table 30: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.
 PEP Chain of Thought Reason for Failure
@@
 "__set_name__")
-Table 27: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.
+Table 31: Cross-Idiom Generalization on JEP & PMD Best Practice Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 31).
+The transfer column indicates training and test data on the left and right side of the arrow.
+Best score across the compared training setups per model are bolded.
 Model Transfer Detection Localization
@@
 Llama3.1-8B-Instruct + SFT + RS-DPO 0.436 0.4898 0.4613 0.0831 0.1351 0.1029
-Table 28: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28).
-The transfer column indicates training and test data on the left and right side of the arrow.
-Best score across the compared training setups per model are bolded.
+Table 32: Egregious Localization Error Types: Breakdown of recurring error categories observed in a manual analysis of 50 egregious localization failures made by the Qwen3-4B non-CoT model trained with SFT+RS-DPO.
+An egregious localization error is defined as predicting a non-empty set of line numbers that is completely disjoint from the non-empty ground-truth set, indicating correct detection but failed localization.
+Error Type Definition Ocurrences (%)
+Adjacent Predicts a line immediately adjacent to the true violating line, but not the correct one. 12%
+Before/After Flags a line that follows the best practice rather than the line that violates it. 4%
+Definition vs. Initialization Flags an initialization instead of the definition that actually constitutes the violation. 4%
+Plausible-Looking Flags a superficially plausible but ultimately incorrect line due to surface-level similarity or insufficient semantic nuance. 46%
+Random Selects an apparently arbitrary line, suggesting a collapse in localization despite correct detection. 28%
+Repetition Repeats the same incorrect line multiple times, indicating a decoding or sampling artifact. 6%